### v1

# Subdiffusive concentration for the chemical distance in Bernoulli percolation

Abstract.
Considering supercritical Bernoulli percolation on ${\mathbb{Z}}^{d}$, Garet and Marchand [GM09] proved a diffusive concentration for the graph distance. In this paper, we sharpen this result by establishing the subdiffusive concentration inequality, which revisits the sublinear bound of the variance proved by Dembin [Dem22] as a consequence. Our approach is inspired by similar work in First-passage percolation [BR08, DHS14], combined with new tools to address the challenge posed by the infinite weight of the model. These tools, including the notion of effective radius and its properties, enable a simple one-step renormalization process as a systematic means of managing the effects of resampling edges.
V.H. Can: Institute of Mathematics, Vietnam Academy of Science and Technology, 18 Hoang Quoc Viet, Cau Giay, Hanoi, Vietnam. cvhao@math.ac.vn
V.Q. Nguyen: Institute of Mathematics, Vietnam Academy of Science and Technology, 18 Hoang Quoc Viet, Cau Giay, Hanoi, Vietnam. nvquyet@math.ac.vn
1. Introduction
1.1. Model and main result
Bernoulli percolation is a simple but well-known probabilistic model for porous material introduced by Broadbent and Hammersley [BH57]. Let $d\geq 2$ and ${{\mathcal{E}}({\mathbb{Z}}^{d})}$ be the set of the edges $e=(x,y)$ of endpoints $x=(x_{1},\ldots,x_{d}),y=(y_{1},\ldots,y_{d})\in{\mathbb{Z}}^{d}$ such that $\|x-y\|_{1}:=\sum_{i=1}^{d}|x_{i}-y_{i}|=1$. Given the parameter $p\in(0,1)$, we let each edge $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ be open with probability $p$ and closed otherwise, independently of the state of other edges. The phase transition of the model has been well-known since 1960s. In particular, it has been shown that there exists a critical parameter $p_{c}(d)\in(0,1)$ such that there is almost surely a unique infinite open cluster ${\mathcal{C}}_{\infty}$ if $p>p_{c}(d)$, whereas all open clusters are finite if $p<p_{c}(d)$. We refer to the book of Grimmett [Gri89] and the reviews of Duminil-Copin [DC18] for classical results and recent developments in percolation theory. In this paper, we are interested in the chemical distance that is the graph distance of infinite cluster in the supercritical regime.
For each $A,B,U\subset{\mathbb{Z}}^{d}$, we define
${\mathrm{D}}^{U}(A,B)$ $:=\inf\{|\gamma|:x\in A,\,y\in B,\,\gamma\text{ is an open nearest-neighbor path from }x\text{ to }y\text{ inside }U\}.$
When $U={\mathbb{Z}}^{d}$, we simply write ${\mathrm{D}}$ for ${\mathrm{D}}^{{\mathbb{Z}}^{d}}$ and we write ${\mathrm{D}}(x,y)$ for ${\mathrm{D}}(\{x\},\{y\})$. Since this graph distance between two disconnected vertices is not well defined, we consider a regularized version as follows. Let $x\in{\mathbb{Z}}^{d}$, we denote by $x^{*}$ the closest point to $x$ in ${\mathcal{C}}_{\infty}$ (in $\|.\|_{\infty}$ distance) and call it the regularized point of $x$. The graph distance is then defined as: for $x,y\in{\mathbb{Z}}^{d}$,
${\mathrm{D}}^{*}(x,y):={\mathrm{D}}(x^{*},y^{*})=\inf\{|\gamma|:\gamma\text{ is an open nearest-neighbor path from }x^{*}\text{ to }y^{*}\}.$
Let ${\mathbf{e}}_{1}=(1,0,\ldots,0)$ and we aim to study the asymptotic behavior of
${\mathrm{D}}^{*}_{n}:={\mathrm{D}}^{*}(0,n{\mathbf{e}}_{1}).$
The linear growth of ${\mathrm{D}}^{*}_{n}$ was described by Garet and Marchand [GM04]: for any $p>p_{c}(d)$, there exists a constant $\mu_{p}({\mathbf{e}}_{1})\in(0,\infty)$ such that,
$\lim_{n\to\infty}\dfrac{{\mathrm{D}}^{*}_{n}}{n}=\mu_{p}({\mathbf{e}}_{1})\quad\text{a.e and in }L^{1}.$ (1.1)
The value $\mu_{p}({\mathbf{e}}_{1})$ is called the time constant. The regularity of the time constant (Lipschitz-continuity of the time constant with respect to $p$) was proved in [CD22] and revisited in [CNN24]. Naturally, the next question we are interested in is the fluctuation and deviation of the graph distance. The large deviation of ${\mathrm{D}}^{*}_{n}$ has been well studied since 1990s, see, for example, [AP96, GM07, DN22]. The moderate deviation of ${\mathrm{D}}^{*}_{n}$ (or precisely the concentration with diffusive scale) was established by Garet and Marchand (see [GM09, Theorem 1.2]), using concentration inequalities of Boucheron, Lugosi, and Massart: for each $C>0$, there exists some constant $c>0$ such that for all $\lambda\in[C\log n,\sqrt{n}$],
$\mathbb{P}[|{\mathrm{D}}^{*}_{n}-{\mathbb{E}}[{\mathrm{D}}^{*}_{n}]|\geq\sqrt{n}\lambda]\leq c^{-1}e^{-c\lambda}.$ (1.2)
Recently, Dembin gave a sublinear bound on the vairance of ${\mathrm{D}}^{*}_{n}$ (see [Dem22, Theorem 1.1]): there exists a positive constant $C$ such that
$\operatorname{Var}({\mathrm{D}}^{*}_{n})\leq\frac{Cn}{\log n}.$ (1.3)
The main result of our paper is to sharpen the moderate deviation (1.2) by establishing a sub-diffusive concentration of ${\mathrm{D}}^{*}_{n}$.
Theorem 1.1.
Let $p>p_{c}(d)$. There exists a positive constant $c>0$ such that for all $n\in{\mathbb{N}}$ and $\kappa\geq 0$,
$\mathbb{P}\left(|{\mathrm{D}}^{*}_{n}-\mathbb{E}[{\mathrm{D}}^{*}_{n}]|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}e^{-c\kappa}.$ (1.4)
Consequently, the sublinear bound of the variance (1.3) holds.
Remark 1.2.
If $\kappa<1$, we can take $c$ small enough such that (1.4) holds trivially. From now on, we focus on the case $\kappa\geq 1$ throughout this paper.
It is worth mentioning that the present work is inspired by similar result in First-passage percolation (FPP) by Damron, Hanson, and Sosoe. We briefly recall here the FPP model and related results. Let $(t_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be i.i.d. random variables with common distribution $\zeta$. Each edge $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ is assigned a random weight $t_{e}$. Then the first passage time is defined in the same way as the chemical distance. Precisely, if $\zeta([0,\infty))=1$, i.e., the edge weights $(t_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ are finite: for $\{x,y\}\subset U\subset{\mathbb{Z}}^{d}$,
${\mathrm{T}}^{U}(x,y):=\inf_{\begin{subarray}{c}\gamma:x\to y\\ \gamma\subset U\end{subarray}}\sum_{e\in\gamma}t_{e},$ (1.5)
where infimum is taken over the set of nearest-neighbor paths inside $U$ from $x$ to $y$. When $U={\mathbb{Z}}^{d}$, we simply write ${\mathrm{T}}$ instead of ${\mathrm{T}}^{{\mathbb{Z}}^{d}}$. If $p_{c}(d)<\zeta([0,\infty))<1$, we set
${\mathrm{T}}^{*}(x,y):={\mathrm{T}}(x^{*},y^{*}),$ (1.6)
where $z^{*}$ denotes the closest point to $z$ in the infinite cluster of finite weight edges. Then the supercritical Bernoulli percolation can be referred as a particular case of FPP with distribution $\zeta=\zeta_{p}=p\delta_{1}+(1-p)\delta_{\infty}$.
The convergence of the scaled passage time in probability to time constant was obtained by Cerf and Théret [CT16, Theorem 4], without any moment assumption: there exists a constant $\mu_{\zeta}(\mathbf{e}_{1})\in[0,\infty)$ such that
$\lim_{n\to\infty}\frac{{\mathrm{T}}^{*}(0,n\mathbf{e}_{1})}{n}=\mu_{\zeta}(\mathbf{e}_{1})\qquad\text{ in probability}.$ (1.7)
In [GM04, Remark 1], Garet and Marchand proved that if ${\mathbb{E}}[t_{e}^{2+\varepsilon}{\mathbb{I}}(t_{e}<\infty)]<\infty$ with some $\varepsilon>0$, then the convergence in (1.7) holds true almost surely and in $L_{1}$. The continuity property of the time constant with respect to $\zeta$ was first shown by Cox and Kesten in [Cox80, CK81, Kes86] for FPP with finite weights. This was later extended for FPP with possibly infinite weights by Garet, Marchand, Procaccia, and Théret in [GMPT17]. It then has been proved by Damron, Hanson, and Sosoe in [DHS15] that if $\zeta([0,\infty))=1$ and ${\mathbb{E}}[t_{e}^{2}\log_{+}t_{e}]<\infty$, then the sublinear variance holds:
$\operatorname{Var}({\mathrm{T}}(0,n{\mathbf{e}}_{1}))\leq Cn/\log n.$ (1.8)
The sublinearity of variance is a particular case of superconcentration, a phenomenon coined by Chatterjee [Cha14] describing the situation that usual concentration tools such as Poicaré or Erfon-Stein inequality give suboptimal bounds. In fact, in some particular models as Gaussian disordered systems [Cha14], he also established a connection between superconcentration and a chaotic phenomenon of the ground states. However, understanding these notions beyond the Gaussian realm remains quite limited. Recently, Ahlberg, Deijfen, and Sfragara [ADS23] showed that this deep relation holds true in the context of First-passage percolation, under the assumption that $\zeta$ is finitely supported and has a finite second moment. Other examples of the superconcentration phenomena can be found in [BDG20, Cha23, DG24].
1.2. Method of the proof
We first remark that in supercritical Bernoulli percolation, the moment conditions of FPP mentioned above do not hold any more, since $\zeta(\{\infty\})=1-p>0$. Hence, we need new tools to handle the infinite weight.
The general strategy of proving the subdiffusive concentration for FPP has been shown in [BR08] and [DHS14]. First, by general bounds, the subdiffusive concentration of ${\mathrm{T}}(0,n{\mathbf{e}}_{1})$ can be reduced to some estimates of the variance of exponential functionals of ${\mathrm{T}}(0,n{\mathbf{e}}_{1})$. Next, to bound the variance of these exponential terms, one could use the entropy inequalities combining with the geometric averaging trick of Benjamini, Kalai, Schramm introduced in their highly influential paper [BKS03]. Moreover, in FPP [DHS14], the key for variance or entropy bounds is to well control for the effect of resampling an edge weight, and the moment conditions play a crucial role in these controls. However, in the context of the graph distance in Bernoulli percolation, closing an edge on the geodesic can have a significant impact on the graph distance due to the possibility of infinite edge-weight values. By introducing the notion effective radius and exploiting the greedy lattice animals, we provide a simple one-step renormalization process as a systematic way to manage the effect of resampling edges. Let us now explain our proof in more detail.
Part A: Subdiffusive concentration for the truncated passage time. Let $(t_{e})_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}$ be a collection of i.i.d. random variables with the same distribution as
$t_{e}=p\delta_{1}+(1-p)\delta_{\log^{2}n}.$
The first passage time ${\mathrm{T}}_{n}$ with truncated weights $(t_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ is defined as
${\mathrm{T}}_{n}={\mathrm{T}}(0,n\mathbf{e}_{1}),$
and our aim is to prove the following.
Theorem 1.3.
There exists a positive constant $c$ such that for all $n\in{\mathbb{N}}$ and $\kappa\geq 0$
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-\mathbb{E}[{{\mathrm{T}}_{n}}]|\geq\sqrt{\tfrac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa).$ (1.9)
Our argument initially follows the common scheme as in [BR08, DHS14]. Considering ${\mathrm{F}}_{m}$ a spatial average of ${\mathrm{T}}_{n}$ defined in (4.3) (inspired by Benjamini, Kalai, and Schramm in [BKS03]), the proof of (1.9) can be transfered to the following. There exists a constant $c>0$ such that with $K=\dfrac{cn}{\log n}$,
$\forall\,|\lambda|<\dfrac{1}{\sqrt{K}},\quad\operatorname{Var}\left[e^{\lambda{{\mathrm{F}}_{m}}}\right]\leq K\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]<\infty.$ (1.10)
Using Falik-Samorodnitsky inequality (Lemma 4.1), the problem (1.10) is then reduced to understand well the behavior of $G=e^{\lambda{\mathrm{F}}_{m}}$ when flipping the state of edges in ${{\mathcal{E}}({\mathbb{Z}}^{d})}$. In fact, we shall show in Section 4 that the above variance bound is closely related to a large deviation estimate
${\mathbb{P}}\Bigg{(}\sum_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}(\nabla_{e}{\mathrm{T}}_{n})^{2}\geq Cn\Bigg{)}\leq\exp(-\sqrt{n}(\log n)^{C}),\qquad\nabla_{e}{\mathrm{T}}_{n}={\mathrm{T}}_{n}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{n}(1,t_{e^{c}}),$ (1.11)
where C is a suitable positive constant, and the discrete derivative $\nabla_{e}{\mathrm{T}}_{n}$ measures the difference of ${\mathrm{T}}_{n}$ when flipping the state of $e$.
In order to get a suitable bound for $\nabla_{e}{\mathrm{T}}_{n}$, we introduce a novel notion called effective radius. Roughly speaking, the effective radius of an edge $e$ is the smallest radius $R_{e}$ ensuring a good path bypassing $e$ inside the annulus ${\mathrm{A}}_{R_{e}}(e)=\Lambda_{3R_{e}}(e)\setminus\Lambda_{R_{e}}(e)$, and hence it measures the effect of the edge $e$ to ${\mathrm{T}}_{n}$ as
$|\nabla_{e}{\mathrm{T}}_{n}|\leq\hat{R}_{e}{\mathbb{I}}(e\in\gamma),\quad\hat{R}_{e}=\min\{C_{*}R_{e},\log^{2}n\},$ (1.12)
with $C_{*}$ being a constant introduced in (2.2) and $\gamma$ is a geodesic of ${\mathrm{T}}_{n}$. While the bound by $C_{*}R_{e}$ follows from using the bypass in the definition of effective radius, the bound by $\log^{2}n$ is straightforward from the definition of the truncated passage time. We also show in Section 2 that the effective radii have two important properties.
  • (a)
    First, they are weakly dependent, in the sense that the event $\{R_{e}\leq t\}$ depends only on the status of edges at a distance at most $C_{*}t$ from $e$.
  • (b)
    Second, they exhibit the finite range exponential decay: ${\mathbb{P}}(R_{e}\geq t)\leq\exp(-ct)$ for all $1\leq t\leq\exp(c\log^{2}n)$ with $c$ a positive constant.
Next, we will need an ingredient, the so-called greedy lattice animals (see Section 3 for more details), to control the total cost $\sum_{e\in\gamma}\hat{R}_{e}^{2}$. Precisely, we decompose
$\sum_{e\in\gamma}\hat{R}_{e}^{2}=\sum_{M=1}^{\log^{2}n}M^{2}\sum_{e\in\gamma}{\mathbb{I}}(\hat{R}_{e}=M),$ (1.13)
and exploit the greedy lattice animals theory (see [DHS14, CN19]) to the Bernoulli random variables $({\mathbb{I}}(\hat{R}_{e}=M))_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}$ with the aid of two properties (a) and (b). In fact, we obtain a much better estimate than (1.11):
${\mathbb{P}}\Bigg{(}\sum_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}(\nabla_{e}{\mathrm{T}}_{n})^{2}\geq Cn\Bigg{)}\leq{\mathbb{P}}\Bigg{(}\sum_{e\in\gamma}\hat{R}_{e}^{2}\geq Cn\Bigg{)}\leq\exp(-n/(\log n)^{C}).$ (1.14)
We emphasize that the two properties (a) and (b) alone are not strong enough to gain the above large deviation estimate, and in fact the truncation by $\log^{2}n$ is crucial for this purpose, as detailed in Lemma 3.3.
Part B: Bound on discrepancy between ${\mathrm{D}}^{*}_{n}$ and ${\mathrm{T}}_{n}$ via effective radii. We aim to prove that $|{\mathrm{D}}^{*}_{n}-{\mathrm{T}}_{n}|$ is negligible with overwhelming probability.
Theorem 1.4.
There exists a positive constant $c$ such that for all $L\geq\log^{2}n$,
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}_{n}|\geq L\right)\leq c^{-1}\exp\left(-c\tfrac{L}{\log L+\log^{2}n}\right).$ (1.15)
The heart of the proof of this theorem is to make a suitable coupling of ${\mathrm{D}}^{*}_{n}$ and ${\mathrm{T}}(0^{*},(n\mathbf{e}_{1})^{*})$. Pick a geodesic of ${\mathrm{T}}(0^{*},(n\mathbf{e}_{1})^{*})$, say $\gamma$. First, we identify $e_{*}$ the closed (or $\log^{2}n$-weight) edge with largest effective radius and find a bypass for it. Hence, we get a new path, say $\eta$, from $0^{*}$ to $(n\textbf{e}_{1})^{*}$ with fewer closed edges than the original path $\gamma$ and $|{\mathrm{T}}(\gamma)-{\mathrm{T}}(\eta)|={\mathcal{O}}(R_{e_{*}})$. Next, apply the same procedure to the new path $\eta$. We continue this covering process until there are no closed edges left. When this process ends, we find a set of closed edges $\Gamma\subset\gamma$ such that
  • (i)
    $\forall e,e^{\prime}\in\Gamma,\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}$,
  • (ii)
    $|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\leq C\sum_{e\in\Gamma}R_{e}$,
with $C$ a positive constant. Hence, our task is now to show
${\mathbb{P}}\Bigg{(}\sum_{e\in\Gamma}R_{e}\geq L\Bigg{)}\leq c^{-1}\exp(-cL/(\log L+\log^{2}n)).$ (1.16)
As mentioned above, achieving a large deviation of the total cost $\sum_{e\in S}R_{e}$ with $S$ a general set, as described in (1.16), is not feasible. In fact, the separable property (i) somehow strengthens the local dependent of $(R_{e})_{e\in\Gamma}$, and enables us to prove this large deviation estimate. We refer to Section 5 for the detailed proof.
Finally, let us highlight the novelty of our method. While the idea of constructing detours that bypass a given edge with reasonable length is natural and frequently applied, previous approaches often relied on multi-step renormalization processes. These processes involved constructing bypasses through a sequence of good boxes at various scales, utilizing a random number of boxes with different but deterministic radii. This sophisticated process is sensitive to the choice of scaling, see in particular [CD22, Proposition 4.5 & Remark 4.6]. In contrast, our method represents a simpler, one-step renormalization that involves a single box with a random radius. This approach allows us to achieve a nearly optimal large deviation estimate (1.14). Intuitively, we encapsulate the multi-scale analysis within the decomposition (1.13) and employ the greedy lattice animals theory to analyze each scale. This approach appears to be robust and can be applied to other problems related to controlling the effect of resampling edges in percolation, such as the Lipschitz-continuity of the time constant, as discussed in [CNN24].
1.3. Proof of the main result assuming Theorems 1.3 and 1.4
Using Theorem 1.4,
${\mathbb{E}}[|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}_{n}|]={\mathcal{O}}(\log^{2}n).$
By the triangle inequality,
$|{\mathrm{D}}^{*}_{n}-{\mathbb{E}}[{{\mathrm{D}}^{*}_{n}}]|\leq|{\mathrm{D}}^{*}_{n}-{{\mathrm{T}}_{n}}|+|{{\mathrm{T}}_{n}}-{\mathbb{E}}[{{\mathrm{T}}_{n}}]|+|{\mathbb{E}}[{{\mathrm{T}}_{n}}]-{\mathbb{E}}[{{\mathrm{D}}^{*}_{n}}]|.$
It follows from the last two estimates that for all $\kappa\geq 1$ and $n$ large enough,
$\mathbb{P}\Big{[}|{{\mathrm{D}}^{*}_{n}}-{\mathbb{E}}[{{\mathrm{D}}^{*}_{n}}]|\geq\kappa\sqrt{\tfrac{n}{\log n}}\Big{]}$ $\leq\mathbb{P}\Big{[}|{\mathrm{T}}_{n}-{\mathbb{E}}[{\mathrm{T}}_{n}]|\geq\tfrac{\kappa}{4}\sqrt{\tfrac{n}{\log n}}\Big{]}+\mathbb{P}\Big{[}|{\mathrm{D}}^{*}_{n}-{\mathrm{T}}_{n}|\geq\tfrac{\kappa}{4}\sqrt{\tfrac{n}{\log n}}\Big{]}.$
By Theorem 1.3,
$\mathbb{P}\Big{[}|{\mathrm{T}}_{n}-{\mathbb{E}}[{\mathrm{T}}_{n}]|\geq\tfrac{\kappa}{4}\sqrt{\tfrac{n}{\log n}}\Big{]}\leq c_{1}^{-1}\exp(-c_{1}\kappa/4),$
for some $c_{1}>0$. In addition, using Theorem 1.4,
$\mathbb{P}\Big{[}|{\mathrm{D}}^{*}_{n}-{\mathrm{T}}_{n}|\geq\tfrac{\kappa}{4}\sqrt{\tfrac{n}{\log n}}\Big{]}$ $\leq c_{2}^{-1}\exp\Big{(}-c_{2}\kappa\tfrac{\sqrt{n}}{(\log n)^{5/2}}\Big{)},$
for some $c_{2}>0$. Finally, combining the last three displays we get Theorem 1.1. ∎
1.4. Organization and notation of this paper
The paper is organized as follows. In Section 2, we present the construction of random effective radius and its application to control the effect of flipping an edge. We study some moments and large deviations of lattice animal of dependent weight in Section 3. In Section 4, we first revisit the concentration inequalities and then prove the subdiffusive concentration of the modified graph distance (Theorem 1.3). Finally, we estimate the discrepancy between the graph distance and its modified version via the covering argument (Theorem 1.4) in Section 5.
To conclude this section, we introduce some notations used in the paper.
  • Integer interval. Given an integer $t\geq 1$, we denote by $[t]:=\{1,2,\ldots,t\}$.
  • Metric. We denote by $\|\cdot\|_{1},\|\cdot\|_{\infty},\|\cdot\|_{2}$ the $l_{1},l_{\infty},l_{2}$ norms, respectively.
  • Box and its boundary. Let $x\in{\mathbb{Z}}^{d}$ and $t>0$, we will denote by $\Lambda_{t}(x):=x+[-t,t]^{d}\cap{\mathbb{Z}}^{d}$ the box centered at $x=(x_{1},\ldots,x_{d})\in{\mathbb{Z}}^{d}$ with radius $t$. For convenience, we briefly write $\Lambda_{t}=[-t,t]^{d}$ for $\Lambda_{t}(0)$. We define the boundary of $\Lambda_{t}(x)$ as $\partial\Lambda_{t}(x):=\Lambda_{t}(x)\setminus\Lambda_{t-1}(x)$.
  • Set distance. For $X,Y\subset{\mathbb{Z}}^{d}$, we denote ${\rm d}_{\infty}(X,Y)$ the distance between $X$ and $Y$ by
    ${\rm d}_{\infty}(X,Y):=\inf\{\|x-y\|_{\infty}:x\in X,y\in Y\}.$
  • Edge distance. For each edge $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, we pick a deterministic rule to represent $e=(x_{e},y_{e})$, for example, $\|x_{e}\|_{1}<\|y_{e}\|_{1}$. For $x\in{\mathbb{Z}}^{d}$ and $e,f\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, we denote by
    $\|e-f\|_{\infty}:=\|x_{e}-x_{f}\|_{\infty}\quad\|e-x\|_{\infty}:=\|x_{e}-x\|_{\infty}.$
  • ${\mathbb{Z}}^{d}$-path and set of ${\mathbb{Z}}^{d}$-paths. For any $\ell$, we say that a sequence $\gamma=(v_{0},\ldots,v_{\ell})$ is a ${\mathbb{Z}}^{d}$-path if for all $i\in[\ell],\|v_{i}-v_{i-1}\|_{1}=1$. The length of $\gamma$ is $\ell$, denoted by $|\gamma|$. For $1\leq i<j\leq\ell$, we denote by $\gamma_{v_{i},v_{j}}$ the sub-path of $\gamma$ from $v_{i}$ to $v_{j}$. In addition, if $v_{i}\neq v_{j}$ for $i\neq j$, then we say that $\gamma$ is self-avoiding. From now on, we will shortly write a path in place of a self-avoiding ${\mathbb{Z}}^{d}$-path. For any path $\gamma$, we denote ${\bf s}(\gamma),{\bf e}(\gamma)$ the starting and ending vertices of $\gamma$, respectively. Given $U\subset{\mathbb{Z}}^{d}$, let $\mathcal{P}(U)$ be the set of all paths in $U$.
  • Open path, open cluster and crossing cluster. Given a Bernoulli percolation on ${\mathbb{Z}}^{d}$ with parameter $p$, let $\mathcal{G}_{p}=({\mathbb{Z}}^{d},\{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}:e\text{ is open}\})$. We say that a path is open if all of its edges are open. An open cluster is a maximal connected component of $\mathcal{G}_{p}$. An open cluster ${\mathcal{C}}$ crosses a box $\Lambda$, if for all $d$ directions, there is an open path in ${\mathcal{C}}\cap\Lambda$ connecting the two opposite faces of $\Lambda$.
[…]
  • Diameter. For $A\subset{\mathbb{Z}}^{d}$ and $1\leq i\leq d$, let us define
    $\text{diam}_{i}(A)=\max_{x,y\in A}|x_{i}-y_{i}|,$
    and we thus denote diam$(A)$ the diameter of $A$ by
    $\text{diam}(A)=\max_{1\leq i\leq d}\text{diam}_{i}(A).$
  • Geodesic. Given $\{x,y\}\subset U\subset{\mathbb{Z}}^{d}$, we denote by $\gamma$ a geodesic between $x$ and $y$ of ${\mathrm{D}}^{U}(x,y)$ if $\gamma$ is an open path inside $U$ such that $|\gamma|={\mathrm{D}}^{U}(x,y)$. We also denote by $\eta$ a geodesic between $x$ and $y$ of ${\mathrm{T}}^{U}(x,y)$ if $\eta$ is a path inside $U$ such that ${\mathrm{T}}(\eta)={\mathrm{T}}^{U}(x,y)$. If there are several choices for $\gamma$ or $\eta$, we choose one of them by a deterministic rule.
2. The effect of resampling
As mentioned in Introduction, the key to the concentration of the truncated passage time ${\mathrm{T}}_{n}$ is understanding the cost of resampling the edge weights along the geodesic of ${\mathrm{T}}_{n}$. To study this issue, given an edge $e$, we introduce the notion of effective radius $R_{e}$, which measures how large the change in passage time is when flipping the weight of $e$. The radius $R_{e}$ indeed guarantees an open bypass of $e$ whose length is comparable to $R_{e}$. Hence, it turns out that the cost of flipping the edge weight is given by (1.12). We then investigate the properties of the effective radii, including the local dependence and light-tailed decay distribution. These properties will play a central role in controlling the total cost of resampling edges.
2.1. Effective radius and its applications
We couple Bernoulli percolation with First-passage percolation where $\zeta=p\delta_{1}+(1-p)\delta_{\log^{2}n}$ as follows: each open edge $e$ has weight $1$ and $\log^{2}n$ otherwise. We now set up some definitions for the notion of effective radius. Recall that ${\mathcal{P}}(A)$ is the set of all paths in $A$. Let us define the set of open paths in $A\subset{\mathbb{Z}}^{d}$ by
$\mathbb{O}{(u,v;A)}:=\{\gamma\in{\mathcal{P}}(A):{\bf s}(\gamma)=u,{\bf e}(\gamma)=v,\textrm{ $\gamma$ is open}\};\quad\mathbb{O}{(A)}:=\bigcup_{u,v\in A}\mathbb{O}{(u,v;A)},$
here recall that ${\bf s}(\gamma)$ and ${\bf e}(\gamma)$ are starting and ending points of $\gamma$. If $A={\mathbb{Z}}^{d}$, we simply write $\mathbb{O}$ and $\mathbb{O}(u,v)$ for $\mathbb{O}({\mathbb{Z}}^{d})$ and $\mathbb{O}(u,v;{\mathbb{Z}}^{d})$, respectively. For $u,v\in A\subset{\mathbb{Z}}^{d}$, we define the set of geodesics of ${\mathrm{T}}^{A}(u,v)$ as
$\mathbb{G}(u,v;A):=\{\gamma\in{\mathcal{P}}(A):{\bf s}(\gamma)=u,{\bf e}(\gamma)=v,{\mathrm{T}}(\gamma)={\mathrm{T}}^{A}(u,v)\},$
and the set of geodesics in $A$ as
$\mathbb{G}(A):=\bigcup_{u,v\in A}\mathbb{G}(u,v;A).$
Also, define the set of modified geodesics by
$\mathbb{G}_{*}(u,v;A)$ ::=:= $\Big{\{}\gamma\in{\mathcal{P}}(A):{\bf s}(\gamma)=u,{\bf e}(\gamma)=v,\exists\,\pi\in\mathbb{G}(A):\gamma\setminus\pi\textrm{ is open}\Big{\}},$
$\mathbb{G}_{*}(A)$ ::=:= $\bigcup_{u,v\in A}\mathbb{G}_{*}(u,v;A),$
with the convention that an empty path is open (particularly, $\mathbb{G}(A),\mathbb{O}(A)\subset\mathbb{G}_{*}(A)$). In other words, $\mathbb{G}_{*}(u,v;A)$ is the set of paths from $u$ to $v$ that lie within $A$, obtained by replacing some segments of a geodesic with open paths. If $A={\mathbb{Z}}^{d}$, we simply write $\mathbb{G}(u,v),\mathbb{G},\mathbb{G}_{*},\mathbb{G}_{*}(u,v)$ for $\mathbb{G}(u,v;{\mathbb{Z}}^{d})$, $\mathbb{G}({\mathbb{Z}}^{d})$, $\mathbb{G}_{*}({\mathbb{Z}}^{d})$, $\mathbb{G}_{*}(u,v;{\mathbb{Z}}^{d})$, respectively.
Remark 2.1.
Given $B\subset A\subset{\mathbb{Z}}^{d}$, if $\gamma\in\mathbb{G}(A)$ (or $\gamma\in\mathbb{G}_{*}(A)$) and $\pi$ is a sub-path of $\gamma$ such that $\pi\subset B$, then $\pi\in\mathbb{G}(B)$ (or $\pi\in\mathbb{G}_{*}(B)$). In addition, $\mathbb{G}(A)$ and $\mathbb{G}_{*}(A)$ are measurable with respect to edge weights inside $A$. Moreover, if $\gamma\in\mathbb{G}_{*}(u,v)$ and $\eta\setminus\gamma$ is open, then $\eta\in\mathbb{G}_{*}(u,v)$ as well.
For each $N\geq 1$, we define the family of $N$-annuli:
$\forall e\in{{\mathcal{E}}({\mathbb{Z}}^{d})},\quad{\mathrm{A}_{N}(e)}:=\Lambda_{3N}(e)\setminus\Lambda_{N}(e),\quad\partial{\mathrm{A}_{N}(e)}:=\partial\Lambda_{3N}(e)\cup\partial\Lambda_{N}(e).$ (2.1)
A path $\eta\subset{\mathrm{A}_{N}(e)}$ is called a crossing path of ${\mathrm{A}_{N}(e)}$ if it joints $\partial\Lambda_{N}(e)$ and $\partial\Lambda_{3N}(e)$. Let ${\mathscr{C}}({\mathrm{A}_{N}(e)})$ be the collection of all crossing paths of ${\mathrm{A}_{N}(e)}$. Suppose that $\gamma$ is a path crossing the annulus ${\mathrm{A}}_{N}(e)$ at least once ($\gamma$ is not necessarily inside ${\mathrm{A}}_{N}(e)$). Then, we denote by $\gamma^{{\bf i},e}\in{\mathscr{C}}({\mathrm{A}_{N}(e)})$ and $\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}_{N}(e)})$ the first and last sub-paths of $\gamma$ crossing ${\mathrm{A}}_{N}(e)$, respectively, ordered from ${\bf s}(\gamma)$ to ${\bf e}(\gamma)$ (see Figure 1-A).
(a) First and last crossing paths
(b) A short open path $\pi_{e}$ joining crossing paths
Figure 1. Illustration of crossing paths and the event $\mathcal{V}^{1}_{e}$
Fix $\rho$ and $c_{*}$ as in Lemmas 2.10 and 2.14, respectively, and define
$C_{*}:=1/c_{*}+(100\rho^{2})^{d};$ (2.2)
and let
$\mathcal{V}_{N}^{1}(e):=\{\forall\,\gamma_{1},\gamma_{2}\in\mathbb{G}_{*}(\Lambda_{C_{*}N}(e))\cap{\mathscr{C}}({\mathrm{A}_{N}(e)}):\,\,{\mathrm{D}}^{{\mathrm{A}_{N}(e)}}(\gamma_{1},\gamma_{2})\leq C_{*}N\};$
$\mathcal{V}_{N}^{2}(e):=\{\forall\,x,y\in\Lambda_{3N}(e)\text{ with }{\mathrm{D}}^{\Lambda_{3N}(e)}(x,y)<\infty:\,\,{\mathrm{D}}^{\Lambda_{4N}(e)}(x,y)\leq C_{*}N\}.$
Definition 2.2.
For each $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, we define the effective radius of $e$ as
$R_{e}:=\inf\{N\geq 1:\mathcal{V}_{N}^{1}(e)\cap\mathcal{V}_{N}^{2}(e)\text{ occurs}\}.$ (2.3)
Roughly speaking, $R_{e}$ is the radius that ensures the existence of an economically feasible bypass for the edge $e$ with a cost comparable to $R_{e}$ (see Figure 1-B). Additionally, as we will see below, $R_{e}$ has beneficial properties such as local dependence and a light-tailed distribution, using the robust connectivity of the supercritical percolation
Remark 2.3.
By the definition of effective radius and Remark 2.1, for any $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ and $t\geq 1$, the event $\{R_{e}=t\}$ depends solely on the state of edges in the box $\Lambda_{C_{*}t}(e)$.
Proposition 2.4.
There exists a constant $c>0$ such that for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ and $1\leq t\leq\exp(c\log^{2}n)$,
$\mathbb{P}(R_{e}\geq t)\leq c^{-1}\exp(-ct).$
Given an edge $e$ in a modified geodesic, the following proposition allows us to build a detour that avoids $e$ with an economical cost that is comparable to the effective radius.
Proposition 2.5.
Let $x,y\in{\mathbb{Z}}^{d}$ and $\gamma\in\mathbb{G}_{*}(x,y)$, and $e$ be an edge in $\gamma$ such that $x,y\notin\Lambda_{3R_{e}}(e)$. Then there exists another path $\eta_{e}\in\mathbb{G}_{*}(x,y)$ such that:
  • (i)
    $\eta_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$, and $\eta_{e}\setminus\gamma\subset{\mathrm{A}}_{R_{e}}(e)$ and is an open path;
  • (ii)
    $|\eta_{e}\setminus\gamma|\leq C_{*}R_{e}$;
Proof.
Since $e\in\gamma$ and $x,y\notin\Lambda_{3R_{e}}(e)$, $\gamma$ crosses the annulus ${\mathrm{A}_{R_{e}}(e)}$ at least twice. Let $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ be these first and last crossing paths of ${\mathrm{A}_{R_{e}}(e)}$, respectively. Then we have $\gamma^{{\bf i},e},\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))$ and $\gamma^{{\bf i},e}\neq\gamma^{{\bf o},e}$. By the hypothesis that $\gamma\in\mathbb{G}_{*}$ and Remark 2.1, both $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ belong to ${\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. Furthermore, by the definition of $R_{e}$, the event $\mathcal{V}^{1}_{R_{e}}(e)$ occurs, and so ${\mathrm{D}}^{{\mathrm{A}_{R_{e}}(e)}}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})\leq C_{*}R_{e}$. Let $\pi_{e}$ be a geodesic of ${\mathrm{D}}^{{\mathrm{A}_{R_{e}}(e)}}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})$. Then $\pi_{e}$ consists of only open edges and satisfies
$|\pi_{e}|={\mathrm{D}}^{{\mathrm{A}_{R_{e}}(e)}}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})\leq C_{*}R_{e}.$
Suppose that $\pi_{e}$ intersects with $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ at $z_{{\bf i}}$ and $z_{{\bf o}}$, respectively. We define
$\eta_{e}:=\gamma_{x,z_{{\bf i}}}\cup\pi_{e}\cup\gamma_{z_{{\bf o}},y}.$
Therefore, $\eta_{e}\setminus\gamma=\pi_{e}$ is an open path. Together with $\gamma\in\mathbb{G}_{*}(x,y)$, we get $\eta_{e}\in\mathbb{G}_{*}(x,y)$, thanks to Remark 2.1. Notice that $|\eta_{e}\setminus\gamma|=|\pi_{e}|\leq C_{*}R_{e}$. Moreover, since $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ are the first and last crossing paths of ${\mathrm{A}}_{R_{e}}(e)$, one has $\gamma_{x,z_{\bf i}}\cap\Lambda_{R_{e}}(e)=\varnothing$ and $\gamma_{z_{\bf o},y}\cap\Lambda_{R_{e}}(e)=\varnothing$. In addition, $\pi_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$ since $\pi_{e}\subset{\mathrm{A}}_{R_{e}}(e)$. Hence, $\eta_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$ and so we obtain the claimed result. ∎
For any path $\gamma$, we denote by $\text{clo}(\gamma)$ the set of all closed edges of $\gamma$. The following result will help us gradually erase closed edges in a geodesic, turning it into an open path with a controllable cost. For more details, see in the proof of Proposition 5.1.
Proposition 2.6.
Let $x,y\in{\mathcal{C}}_{\infty}$. Let $\gamma\in\mathbb{G}_{*}(x,y)$, and let $e$ be an edge in ${\mathrm{clo}}(\gamma)$ such that $\{x,y\}\not\subset\Lambda_{3R_{e}}(e)$. Then there exists another path $\eta_{e}\in\mathbb{G}_{*}(x,y)$ such that:
  • (i)
    ${\mathrm{clo}}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing$, and $\eta_{e}\setminus\gamma\subset\Lambda_{4R_{e}}(e)$ is an open path;
  • (ii)
    $|\eta_{e}\setminus\gamma|\leq 2C_{*}R_{e}$.
Proof.
Let $\gamma\in\mathbb{G}_{*}(x,y)$ with $x,y\in{\mathcal{C}}_{\infty}$ and let $e\in\text{clo}(\gamma)$ satisfying $\{x,y\}\not\subset\Lambda_{3R_{e}}(e)$. We consider three cases:
Case 1: $x,y\notin\Lambda_{3R_{e}}(e)$. Then, by Proposition 2.5, there exists another path $\eta_{e}\in\mathbb{G}_{*}(x,y)$, such that the path $\eta_{e}\setminus\gamma$ is open, $\eta_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$, and $|\eta_{e}\setminus\gamma|\leq C_{*}R_{e}$. Hence, we get the desired result.
Case 2: $x\in\Lambda_{3R_{e}}(e)$ and $y\notin\Lambda_{3R_{e}}(e)$. The path $\gamma$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once. We call the last crossing path by $\gamma^{{\bf o},e}$. Using $\gamma\in\mathbb{G}_{*}$ and Remark 2.1, we get $\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. As $x\in{\mathcal{C}}_{\infty}$, there exists an open path $\xi_{x,\infty}$ joining $x$ to $\infty$.
Case 2a: $x\in\Lambda_{R_{e}}(e)$. In this case, the open path $\xi_{x,\infty}$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once. Let $\gamma^{{\bf i},e}\subset\xi_{x,\infty}$ be the first crossing path of ${\mathrm{A}_{R_{e}}(e)}$, so $\gamma^{{\bf i},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. Since the event ${\mathcal{V}}^{1}_{R_{e}}(e)$ occurs, there exists an open path $\tilde{\eta}_{e}$, a geodesic of ${\rm D}^{{\mathrm{A}}_{R_{e}}(e)}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})$, satisfying $|\tilde{\eta}_{e}|\leq C_{*}R_{e}$. Suppose that $\tilde{\eta}_{e}$ intersects with $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ at $z_{\bf i}$ and $z_{\bf o}$, respectively (see Figure 2 for illustration). By the definition of $\gamma^{{\bf i},e}$ and $z_{\bf i}$, we have that $\xi_{x,z_{\bf i}}$, the subpath of $\xi_{x,\infty}$ from $x$ to $z_{\bf i}$, is open and satisfies $\xi_{x,z_{\bf i}}\subset\Lambda_{3R_{e}}(e)$. Thus, ${\mathrm{D}}^{\Lambda_{3R_{e}}(e)}(x,z_{\bf i})<\infty$. Thanks to the definition of $\mathcal{V}^{2}_{R_{e}}(e)$, ${\mathrm{D}}^{\Lambda_{4R_{e}}(e)}(x,z_{\bf i})\leq C_{*}R_{e}$. Let us denote by $\tilde{\xi}_{x,z_{\bf i}}$ the geodesic of ${\mathrm{D}}^{\Lambda_{4R_{e}}(e)}(x,z_{\bf i})$ and define
$\eta_{e}:=\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\cup\gamma_{z_{\bf o},y}.$
Figure 2. Case 2a: $x\in\Lambda_{R_{e}}(e)$
We observe that the subpath $\eta_{e}\setminus\gamma\subset\{\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\}$ consists of only open edges, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. Moreover,
$|\eta_{e}\setminus\gamma|\leq|\tilde{\xi}_{x,z_{\bf i}}|+|\tilde{\eta}_{e}|\leq 2C_{*}R_{e},\quad\text{clo}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing,$
since $\{\eta_{e}\cap\Lambda_{R_{e}}\}\subset\tilde{\xi}_{x,z_{\bf i}}$ is open. Hence, the result follows.
Case 2b: $x\in{\mathrm{A}}_{R_{e}}(e)$. Let $\gamma_{x,e}:=\gamma_{x,x_{e}}$ be the subpath of $\gamma$ joining $x$ to $e$. Let $\tilde{x}$ be the first point where $\xi_{x,\infty}$ touches $\partial\Lambda_{3R_{e}}(e)$, and so $\xi_{\tilde{x},x}\subset\Lambda_{3R_{e}}(e)$. Then, $\eta_{\tilde{x},e}:=\xi_{\tilde{x},x}\cup\gamma_{x,e}$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once as $\xi_{\tilde{x},x}\cap\partial\Lambda_{3R_{e}}(e)\neq\varnothing$ and $\gamma_{x,e}\cap\partial\Lambda_{R_{e}}(e)\neq\varnothing$. Notice that since $\gamma\in\mathbb{G}_{*}$, we have $\gamma_{e,x}\in\mathbb{G}_{*}$, and so $\eta_{\tilde{x},e}\in\mathbb{G}_{*}$. Let $\gamma^{\textbf{i},e}\subset\eta_{\tilde{x},e}$ be the first crossing path of ${\mathrm{A}_{R_{e}}(e)}$. Thanks to Remark 2.1 again, $\gamma^{\textbf{i},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. By the definition of $\mathcal{V}^{1}_{R_{e}}(e)$, there exists a geodesic of ${\rm D}^{{\mathrm{A}}_{R_{e}}(e)}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})$ inside ${\mathrm{A}}_{R_{e}}(e)$, denoted by $\tilde{\eta}_{e}$, that includes only open edges and satisfies $|\tilde{\eta}_{e}|\leq C_{*}R_{e}$. Suppose that $\tilde{\eta}_{e}$ intersects with $\gamma^{\textbf{i},e}$ and $\gamma^{\textbf{o},e}$ at $z_{\bf i}$ and $z_{\bf o}$, respectively.
If $z_{\bf i}\in\tilde{\eta}_{e}\cap\gamma_{x,e}$ (see Figure 3-A for illustration), we set
$\eta_{e}:=\gamma_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\cup\gamma_{z_{\bf o},y}.$
(a) Subcase 2b: $z_{\bf i}\in\tilde{\eta}_{e}\cap\gamma_{x,e}$
(b) Subcase 2b: $z_{\bf i}\in\tilde{\eta}_{e}\cap\xi_{\tilde{x},x}$
Figure 3. Case 2b: $x\in{\mathrm{A}}_{R_{e}}(e)$
Therefore, the path $\eta_{e}\setminus\gamma=\tilde{\eta}_{e}$ is open, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. In addition,
$|\eta_{e}\setminus\gamma|=|\tilde{\eta}_{e}|\leq C_{*}R_{e},\quad\eta_{e}\cap\Lambda_{R_{e}}(e)=\varnothing.$
If $z_{\bf i}\in\tilde{\eta}_{e}\cap\xi_{\tilde{x},x}$ (see Figure 3-B for illustration), then ${\mathrm{D}}^{\Lambda_{3R_{e}}(e)}(x,z_{\bf i})<\infty$. By the definition of $\mathcal{V}^{2}_{R_{e}}(e)$, we get that ${\mathrm{D}}^{\Lambda_{4R_{e}}(e)}(x,z_{\bf i})\leq C_{*}R_{e}$. We take $\tilde{\xi}_{x,z_{\bf i}}$ a geodesic of ${\mathrm{D}}^{\Lambda_{4R_{e}}(e)}(x,z_{\bf i})$ and set
$\eta_{e}:=\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\cup\gamma_{z_{\bf o},y}.$
Hence, $\tilde{\eta}_{e}\subset\eta_{e}\setminus\gamma\subset\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}$ consists of only open edges, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. Also,
$|\eta_{e}\setminus\gamma|\leq|\tilde{\xi}_{x,z_{\bf i}}|+|\tilde{\eta}_{e}|\leq 2C_{*}R_{e},\quad\text{clo}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing,$
as $\{\eta_{e}\cap\Lambda_{R_{e}}\}\subset\tilde{\xi}_{x,z_{\bf i}}$ is open. We complete the proof for this case.
Case 3: $x\notin\Lambda_{3R_{e}}(e)$ and $y\in\Lambda_{3R_{e}}(e)$. The proof is similar to that of Case 2 and is omitted. ∎
Next, we provide an estimate for the impact of resampling an edge on the first passage time via a truncation of effective radius.
Proposition 2.7.
Let $(\hat{R}_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be the sequence of truncated effective radii defined as
$\forall\,e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}:\quad\hat{R}_{e}:=\min\{C_{*}R_{e},\,\log^{2}n\}.$ (2.4)
Then for any $x,y\in{\mathbb{Z}}^{d}$ and $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$,
$0\leq{\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})\leq(\log^{2}n{\mathbb{I}}({\mathcal{U}}_{e})+\hat{R}_{e}){\mathbb{I}}(e\in{\gamma}),$ (2.5)
where ${\mathrm{T}}_{x,y}(t_{e},t_{e^{c}})$ denotes ${\mathrm{T}}(x,y)$ as a function of weights $t_{e}$ and $(t_{e^{\prime}})_{e^{\prime}\neq e}$, and $\gamma$ is a geodesic of ${\mathrm{T}}_{x,y}(1,t_{e^{c}})$, and
${\mathcal{U}}_{e}:=\{3R_{e}\geq\min\{\|e-x\|_{\infty},\|e-y\|_{\infty}\}\}.$
Proof.
Since ${\mathrm{T}}_{x,y}$ is increasing in $t_{e}$, the first inequality in (2.5) is trivial. On the other hand, if $e\notin\gamma$ then
${\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})=0.$
It is obvious that
${\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})\leq\log^{2}n.$
Suppose that $e\in\gamma$ and ${\mathcal{U}}_{e}^{c}$ occurs (and thus neither $x$ nor $y$ is in $\Lambda_{3R_{e}}(e)$). Applying Proposition 2.5 to $\gamma\in\mathbb{G}_{*}(x,y)$ and $e\in\gamma$, there exists a path $\eta_{e}$ between $x$ and $y$ such that
${\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})\leq{\mathrm{T}}(\eta_{e}\setminus\gamma_{z})=|\eta_{e}\setminus\gamma_{z}|\leq C_{*}R_{e}.$
Combining the last three estimates, we get the desired result. ∎
2.2. Proof of Proposition 2.4
We first review some intrinsic properties of chemical distance and crossing cluster of supercritical Bernoulli percolation.
Lemma 2.8.
[Pis96, Theorem 2] There exists a constant $c>0$ such that for all $x\in{\mathbb{Z}}^{d}$ and $t>0$
$\mathbb{P}(\|x-x^{*}\|_{\infty}\geq t)\leq c^{-1}\exp(-ct^{d-1}).$ (2.6)
Lemma 2.9.
[Gri89, Theorem 7.68] There exists a constant $c>0$ such that for all $t>0$,
$\mathbb{P}(\Lambda_{t}\text{ has an }\text{open crossing cluster})\geq 1-c^{-1}\exp(-ct).$
Lemma 2.10.
[AP96, (4.49)] There exist $\rho,c>0$ such that for every $x\in{\mathbb{Z}}^{d}$ and $t\geq\rho\|x\|_{\infty}$,
$\max\{\mathbb{P}(\infty>{\mathrm{D}}(0,x)\geq t),\mathbb{P}({\mathrm{D}}^{*}(0,x)\geq t)\}\leq c^{-1}\exp(-ct).$ (2.7)
Lemma 2.11.
[Gri89, Lemma 7.104] For any $\varepsilon>0$, there exists $c>0$ such that for all $t>0$,
(two disjoint open clusters with diameter at least εt in $\mathbb{P}(\,\exists\,\text{two disjoint open clusters with diameter at least $\varepsilon t$ in $\Lambda_{t}$})\leq c^{-1}\exp(-ct).$
2.2.1. Good box
We present in this part a condition called good box under which the bypass in the definition of effective radius can be easily constructed. For $m\in{\mathbb{N}}$ and $X\subset{\mathbb{Z}}^{d}$, let ${\mathcal{B}}_{m}(X)$ denote the set of all boxes of side-length $m$ in $X$. Let $\rho$ be the constant defined as in Lemma 2.10, and set
$N_{\rho}:=\lfloor\tfrac{N}{16\rho^{2}}\rfloor.$
Roughly speaking, a good box possesses the large-scale geometry of its percolation cluster so similar to Euclidean space that guarantees the feasibility of constructing the effective radius.
Definition 2.12.
For each $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, we say that the box $\Lambda_{3N}(e)$ is good if the following hold:
  • (i)
    there exists an open cluster ${\mathcal{C}}$ in $\Lambda_{3N}(e)$ that contains a crossing cluster of $\Lambda$ for all $\Lambda\in{\mathcal{B}}_{N_{\rho}}(\Lambda_{3N}(e))$;
  • (ii)
    for all $x,y\in\Lambda_{3N}(e)$ with $\|x-y\|_{\infty}\leq 2N_{\rho}$, if ${\mathrm{D}}(x,y)<\infty$, then ${\mathrm{D}}^{\Lambda_{4N}(e)}(x,y)={\mathrm{D}}(x,y)\leq 4\rho N_{\rho}$;
  • (iii)
    for all $x,y\in{\mathrm{A}_{N}(e)}$ with ${\rm d}_{\infty}(\{x,y\},\partial{\mathrm{A}_{N}(e)})\geq N/2$ and $\|x-y\|_{\infty}\leq 2N_{\rho}$, if ${\mathrm{D}}(x,y)<\infty$, then ${\mathrm{D}}^{{\mathrm{A}_{N}(e)}}(x,y)={\mathrm{D}}(x,y)\leq 4\rho N_{\rho}$;
  • (iv)
    if $\pi\in{\mathcal{P}}(\Lambda_{3N}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N}(e))$ satisfies $\operatorname{Diam}(\pi)\geq N_{\rho}$, then $\pi\cap{\mathcal{C}}\neq\varnothing$.
Lemma 2.13.
There exists a constant $c>0$ such that for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ and $1\leq N\leq\exp(c\log^{2}n)$,
$\mathbb{P}(\Lambda_{3N}(e)\text{ is }\textbf{good})\geq 1-c^{-1}\exp(-cN).$
Among the properties of a good box, (iv) is highly non-trivial and is treated in the following result.
Lemma 2.14.
For any $p>p_{c}(d)$, there exists a positive constant $c_{*}\in(0,1)$ such that for all $t\geq 1/c_{*}$ and $1\leq N\leq\exp(c_{*}\log^{2}n)$,
$\mathbb{P}({\mathcal{E}}_{N})\leq c_{*}^{-1}\exp(-c_{*}N),$
where
${\mathcal{E}}_{N}:=\{\exists\,\textrm{crossing cluster }{\mathcal{C}}\subset\Lambda_{2N},\,\exists\,\pi\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(\Lambda_{tN}),$ $\operatorname{Diam}(\pi)\geq N/2:\,\pi\cap{\mathcal{C}}=\varnothing\}.$
Proof.
By scaling $c_{*}$, it is sufficient to prove the result for $N$ sufficiently large. Hence, in the following computations, we assume that $N$ is large enough. In the case $d=2$, by the standard arguments using the planarity of ${\mathbb{Z}}^{2}$, we can prove the claim even for general paths $\pi$. Now we assume that $d\geq 3$. Let us define
${\mathcal{L}}_{N}:=\{\forall\text{ crossing cluster }{\mathcal{C}}\subset\Lambda_{2N},\,\forall\text{ open cluster }{\mathcal{D}}\subset\Lambda_{2N}\text{ with }\operatorname{Diam}({\mathcal{D}})\geq N:{\mathcal{C}}\cap{\mathcal{D}}\neq\varnothing\}.$
Then we have
$({\mathcal{E}}^{\prime}_{N})^{c}\cap{\mathcal{L}}_{N}\subset({\mathcal{E}}_{N})^{c},$ (2.8)
where
${\mathcal{E}}^{\prime}_{N}:=\{\exists\,\pi\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(\Lambda_{tN}),\operatorname{Diam}(\pi)\geq N/2:\pi\not\longleftrightarrow\partial\Lambda_{2N}\}.$
By Lemma 2.11, there exists $c_{1}>0$ such that ${\mathbb{P}}({\mathcal{L}}_{N}^{c})\leq c_{1}^{-1}\exp(-c_{1}n)$, and thus (2.8) gives
$\mathbb{P}({\mathcal{E}}_{N})\leq\mathbb{P}({\mathcal{E}}^{\prime}_{N})+c_{1}^{-1}\exp(-c_{1}N).$ (2.9)
Suppose that ${\mathcal{E}}^{\prime}_{N}$ occurs, and let $\pi\in{\mathcal{P}}(\Lambda_{N})\cap\mathbb{G}_{*}(\Lambda_{tN})$ be a path that realizes ${\mathcal{E}}^{\prime}_{N}$. We assume further that the diameter of $\pi$ is achieved in the first coordinate direction, i.e., there exist $u,v\in\pi$ with $u_{1}-v_{1}=\operatorname{Diam}(\pi)=N/2$. Let $\pi^{*}$ be the subpath of $\pi$ from $u$ to $v$. Since $\pi\in\mathbb{G}_{*}(\Lambda_{tN})$, so is $\pi^{*}$.
We claim that for any $a>0$, there exists $\varepsilon\in(0,1/3)$ such that for all $N\leq\exp(\varepsilon\log^{2}n)$ and $t\geq 1/\varepsilon$,
${\mathbb{P}}(({\mathcal{L}}_{N}^{\prime})^{c})\leq\varepsilon^{-1}\exp(-\varepsilon N),\textrm{ where }{\mathcal{L}}_{N}^{\prime}:=\{\forall\,\gamma\in\cup_{x,y\in\Lambda_{N}}\mathbb{G}(x,y;\Lambda_{tN}):|{\mathrm{clo}}(\gamma)|\leq\tfrac{aN}{\log N}\},$ (2.10)
where recall that ${\mathrm{clo}}(\gamma)=\{e\in\gamma:e\textrm{ is closed}\}$. Indeed, by Lemmas 2.8 and 2.10 with the remark that $d\geq 3$, there exists $\alpha\in(0,1)$ such that
$\mathbb{P}({\mathcal{A}}_{N})\geq 1-\alpha^{-1}\exp(-\alpha N),$
where
${\mathcal{A}}_{N}:=\{\forall x\in\Lambda_{N}:\|x-x^{*}\|_{1}\leq\sqrt{N}/2\}\cap\{\forall x,y\in\Lambda_{2N}\cap{\mathcal{C}}_{\infty}:{\mathrm{D}}(x,y)\leq\alpha^{-1}N\}.$
On ${\mathcal{A}}_{N}$, for all $t\geq 2/\alpha$ and $x,y\in\Lambda_{N}$,
${\mathrm{T}}^{\Lambda_{tN}}(x,y)$ $\leq{\mathrm{T}}^{\Lambda_{tN}}(x,x^{*})+{\mathrm{T}}^{\Lambda_{tN}}(x^{*},y^{*})+{\mathrm{T}}^{\Lambda_{tN}}(y^{*},y)$
$\leq\log^{2}n(\|x-x^{*}\|_{1}+\|y-y^{*}\|_{1})+{\mathrm{D}}(x^{*},y^{*})\leq\sqrt{N}\log^{2}n+\alpha^{-1}N.$
Therefore, for all $\gamma\in\cup_{x,y\in\Lambda_{N}}\mathbb{G}(x,y;\Lambda_{tN})$,
$|{\mathrm{clo}}(\gamma)|\leq\frac{{\mathrm{T}}^{\Lambda_{tN}}(x,y)}{\log^{2}n}\leq\sqrt{N}+\frac{\alpha^{-1}N}{\log^{2}n}\leq aN/\log N,$
for all $1/\varepsilon\leq N\leq\exp(\varepsilon\log^{2}n)$ with $\varepsilon=\varepsilon(a,\alpha)$ sufficiently small. Hence, (2.10) follows.
Let $L\asymp\log N$ be an integer chosen later and define $K:=\lfloor\tfrac{N}{2L}\rfloor$. For $0\leq q\leq K$, we define surfaces and slabs:
$H_{u,q}:=\{x=(x_{1},\ldots,x_{d})\in\Lambda_{2N}:x_{1}=u_{1}+qL\};$
$S_{u,q}:=\{x=(x_{1},\ldots,x_{d})\in\Lambda_{2N}:u_{1}+(q-1)L\leq x_{1}<u_{1}+qL\},$
where recall that $u$ is the starting point of $\pi^{*}$. Notice that for all $q\in[K]$, $\pi^{*}$ has to cross the slab $S_{u,q}$ from $H_{u,q-1}$ to $H_{u,q}$. For each $q\in[K]$, let $\pi^{*}_{q}\subset\pi^{*}$ denote the connected subpath of $\pi^{*}\cap S_{u,q}$ such that $H_{u,q-1}\xleftrightarrow{\Lambda_{N}}H_{u,q}$ via $\pi^{*}_{q}$. Hence, on the event ${\mathcal{E}}^{\prime}_{N}$,
[…]
Let $I:=\{q\in[K]:\pi_{q}^{*}\in{\mathbb{O}}(\Lambda_{N})\}$ be the set of open subpaths of $\pi^{*}$. As $\pi^{*}\in\mathbb{G}_{*}(u,v;\Lambda_{tN})$, there exists $\eta\in\mathbb{G}(u,v;\Lambda_{tN})$ such that $\pi^{*}\setminus\eta$ is open. Then for any $a>0$, conditioned on ${\mathcal{L}}_{N}^{\prime}$ (defined in (2.10) with $t\geq 1/\varepsilon$ and $\varepsilon=\varepsilon(a)$ sufficiently small), we get that
$\textrm{clo}(\pi^{*})\leq\textrm{clo}(\eta)\leq\frac{aN}{\log N},$
which implies that
$|I|\geq K-|{\mathrm{clo}}(\pi^{*})|\geq K-\frac{aN}{\log N}\geq\frac{N}{4L},$ (2.11)
when $L\leq\log N/(5a)$. Therefore,
$\mathbb{P}({\mathcal{E}}^{\prime}_{n}\cap{\mathcal{L}}^{\prime}_{N})$
[…]
where in the last line we have used the independence of events in disjoint slabs. By Lemmas 2.9, 2.11 and union bound, there exists a positive constant $C$ such that if $L\geq C\log N$,
[…]
Thus, for all $N$ sufficiently large
$\mathbb{P}({\mathcal{E}}^{\prime}_{n}\cap{\mathcal{L}}^{\prime}_{N})\leq\exp(-N/(8C)).$ (2.12)
Finally, we clarify the choice of parameters. We take $a=1/(10C)$ and $L=\lceil C\log N\rceil$ under which $L\leq\log N/(5a)$ and $L\geq C\log N$. Using (2.12) and (2.10),
$\mathbb{P}({\mathcal{E}}^{\prime}_{n})\leq\varepsilon^{-1}\exp(-\varepsilon N)+\exp(-N/(8C)),$
with $\varepsilon=\varepsilon(a)$ given in (2.10). Combining this with (2.9) and taking $c_{*}=\frac{1}{4}\min\{\varepsilon,c_{1},1/8C\}$, we obtain the desired result. ∎
Proof of Lemma 2.13.
Fix $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. It follows from Lemmas 2.9 and 2.11 that
$\mathbb{P}(\Lambda_{3N}(e)\text{ does not satisfy (i)})\leq c_{1}^{-1}\exp(-c_{1}N_{\rho}),$ (2.13)
for some constant $c_{1}>0$. Next, by Lemma 2.10, there exists $c_{2}>0$ such that
[…]
$\leq\mathbb{P}(\exists\,x,y\in\Lambda_{3N}(e),\|x-y\|_{\infty}<2N_{\rho},{\mathrm{D}}(x,y)<\infty,{\mathrm{D}}(x,y)\geq 4\rho N_{\rho})\leq c_{2}^{-1}\exp(-c_{2}N).$ (2.14)
To deal with (iii), we observe that if the property (iii) is not satisfied, then
(iii’) $\exists\,x,y\in{\mathrm{A}_{N}(e)}\text{ such that }{\rm d}_{\infty}(\{x,y\},\partial{\mathrm{A}_{N}(e)})\geq N/2,\|x-y\|_{\infty}\leq 2N_{\rho},{\mathrm{D}}(x,y)\in[4\rho N_{\rho},\infty).$
Thanks to union bound and Lemma 2.10 again, there exists $c_{3}>0$ such that
[…]
$\leq c_{3}^{-1}|{\mathrm{A}_{N}(e)}|^{2}\exp(-c_{3}N_{\rho})\leq c_{3}^{-1}\exp(-c_{3}^{2}N_{\rho}).$ (2.15)
Suppose now that $\Lambda_{3N}(e)$ satisfies (i) but not (iv). Then it holds that
[…]
Notice that given a path $\pi\subset\Lambda_{3N}(e)$ with $\operatorname{Diam}(\pi)\geq N_{\rho}$, there exist $x,y\in\pi$ such that $\|x-y\|_{\infty}\geq N_{\rho}$. That is, there exists a sub-path $\pi^{\prime}\subset\pi_{x,y}\cap\Lambda_{N_{\rho}/4}(x)$ such that $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$, where $\pi_{x,y}$ is the sub-path of $\pi$ from $x$ to $y$. Notice that we always can find $\Lambda\in\mathcal{B}_{N_{\rho}/2}(\Lambda_{3N}(e))$ such that $\Lambda$ contains $\Lambda_{N_{\rho}/4}(x)\cap\Lambda_{3N}(e)$, and hence $\pi^{\prime}\subset\Lambda$ with $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$. Therefore, it follows from (iv’) that there exists a vertex $z\in\Lambda_{3N}(e)$, a crossing cluster ${\mathcal{C}}^{\prime}\subset\Lambda_{N_{\rho}/2}(z)\subset\Lambda_{3N}(e)$, and a sub-path $\pi^{\prime}\in{\mathcal{P}}(\Lambda_{N_{\rho}/4}(z))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N_{\rho}/4}(z))$ with $\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8$, such that $\pi^{\prime}\cap{\mathcal{C}}^{\prime}=\varnothing$. Using Lemma 2.14, there exists $c_{4}>0$ such that for all $N\geq 1$
[…]
$\leq\mathbb{P}\Big{(}\exists z\in\Lambda_{3N}(e),\exists\text{ a crossing cluster }{\mathcal{C}}^{\prime}\subset\Lambda_{N_{\rho}/2}(z),$
$\qquad\qquad\qquad\qquad\qquad\qquad\exists\,\pi^{\prime}\in{\mathcal{P}}(\Lambda_{N_{\rho}/4}(x))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N_{\rho}/4}(x)),\operatorname{Diam}(\pi^{\prime})\geq N_{\rho}/8:\pi^{\prime}\cap{\mathcal{C}}^{\prime}=\varnothing\Big{)}$
$\leq c_{4}^{-1}N^{d}\exp(-c_{4}N_{\rho}).$
Combining this estimate with (2.13), (2.2.1), and (2.2.1) yields the desired result. ∎
2.2.2. Proof of Proposition 2.4
Fix $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. Using the definition of $R_{e}$, we have for each $t\geq 2$,
$\mathbb{P}(R_{e}\geq t)\leq 1-\mathbb{P}(\mathcal{V}_{t-1}^{1}(e)\cap\mathcal{V}_{t-1}^{2}(e)).$
Using this estimate and Lemma 2.13, the result follows once we can prove
$\{\Lambda_{3N}(e)\text{ is }\textbf{ good}\}\subset\mathcal{V}^{1}_{N}(e)\cap\mathcal{V}^{2}_{N}(e).$ (2.16)
To this end, we assume that $\Lambda_{3N}(e)\text{ is }\textbf{good}$. We first prove that $\mathcal{V}^{1}_{N}(e)$ holds. Let $\gamma_{1},\gamma_{2}\in\mathbb{G}_{*}(\Lambda_{C_{*}N}(e))\cap{\mathscr{C}}({\mathrm{A}_{N}(e)})$. For each $j\in\{1,2\}$, there exists a connected subpath $\pi_{j}\subset\gamma_{j}\cap\left\{\Lambda_{2N+\frac{N_{\rho}}{2}}(e)\setminus\Lambda_{2N-\frac{N_{\rho}}{2}}(e)\right\}$ satisfying
$\forall j\in\{1,2\}\quad\pi_{j}\in{\mathcal{P}}(\Lambda_{3N}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}N}(e)),\quad\operatorname{Diam}(\pi_{j})\geq N_{\rho},\quad\text{d}_{\infty}(\pi_{j},\partial{\mathrm{A}_{N}(e)})\geq 3N/4.$
Then by Definition 2.12 (iv), we have $\pi_{1}\cap{\mathcal{C}}\neq\varnothing$ and $\pi_{2}\cap{\mathcal{C}}\neq\varnothing$, with ${\mathcal{C}}$ the cluster crossing all sub-boxes of side-length $N_{\rho}$ of $\Lambda_{3N}(e)$. Therefore, there exist $u,v\in{\mathrm{A}_{N}(e)}$ such that $u\in\pi_{1}\cap{\mathcal{C}}$, $v\in\pi_{2}\cap{\mathcal{C}}$, and ${\rm d}_{\infty}(\{u,v\},\partial{\mathrm{A}_{N}(e)})\geq 3N/4$. Moreover, since ${\mathcal{C}}$ contains a crossing cluster in $\Lambda$ for all $\Lambda\in\mathcal{B}_{N_{\rho}}(\Lambda_{3N}(e))$, we can find a sequence of vertices $(x_{i})_{i=0}^{h}\subset{\mathcal{C}}$ with $h\leq(6N/N_{\rho})^{d}\leq(100\rho^{2})^{d}$ such that
$x_{0}=u,\,\,x_{h}=v,\quad{\mathrm{d}}_{\infty}(x_{i},\partial{\mathrm{A}_{N}(e)})\geq N/2,\quad\|x_{i-1}-x_{i}\|_{\infty}\leq 2N_{\rho}\quad\forall\,i\in[h].$
Remark further that ${\mathrm{D}}(x_{i-1},x_{i})<\infty$, as $(x_{i})_{i=0}^{h}\subset{\mathcal{C}}$. Hence, it follows from Definition 2.12 (iii) that ${\mathrm{D}}^{{\mathrm{A}_{N}(e)}}(x_{i-1},x_{i})\leq 4\rho N_{\rho}\leq N/4$. Therefore, $\mathcal{V}^{1}_{N}(e)$ holds since
${\rm D}^{{\mathrm{A}_{N}(e)}}(\gamma_{1},\gamma_{2})\leq\sum_{i=1}^{h}{\mathrm{D}}^{{\mathrm{A}_{N}(e)}}(x_{i-1},x_{i})\leq(6N/N_{\rho})^{d}(4\rho N_{\rho})\leq C_{*}N.$
Next, we assert that $\mathcal{V}^{2}_{N}(e)$ holds. For $x,y\in\Lambda_{3N}(e)$ with ${\mathrm{D}}^{\Lambda_{3N}(e)}(x,y)<\infty$, there exists an open path $\gamma\subset\Lambda_{3N}(e)$ joining $x$ to $y$. Thus, we can find a sequence $(z_{i})_{i=0}^{h}\subset\gamma$ with $h\leq(6N/N_{\rho})\leq(100\rho^{2})^{d}$ such that $z_{0}=x,z_{h}=y$ and
$\forall i\in[h],\quad\|z_{i-1}-z_{i}\|_{\infty}\leq 2N_{\rho}.$ (2.17)
Then, by Definition 2.12 (ii), ${\mathrm{D}}^{\Lambda_{4N}(e)}(y_{i-1},y_{i})\leq 4\rho N_{\rho}$. Therefore,
${\mathrm{D}}^{\Lambda_{4N}(e)}(x,y)\leq\sum_{i=1}^{h}{\mathrm{D}}^{\Lambda_{4N}(e)}(y_{i-1},y_{i})\leq h(4\rho N_{\rho})\leq C_{*}N,$
and thus $\mathcal{V}^{2}_{N}(e)$ holds. In summary, we have proved (2.16) and completed the proof of this lemma. ∎
3. Lattice animals of dependent weight
In the previous section, we derive that the impact of resampling an edge along a modified geodesic can be controlled by its effective radius. Our current goal is to estimate the total cost of rampling, i.e., the sum of the effective radii along a random set. Fortunately, although these effective radii are not independent, they exhibit relatively local dependence (Remark 2.3) and furthermore, satisfy the exponentially decaying tail probability (Proposition 2.4). Moreover, the theory of greedy lattice animals can help us control the total weight of paths in locally dependent environments.
Given an integer $M\geq 1$ and positive constants $a,A$, suppose that $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ is a collection of Bernoulli random variables satisfying
  • (E1)
    $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ are $aM-$dependent, i.e. for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, the variable $I_{e,M}$ is independent of all variables $(I_{e^{\prime},M})_{e^{\prime}\not\in\Lambda_{aM}(e)}$.
  • (E2)
    $q_{M}:=\sup_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}{\mathbb{E}}[I_{e,M}]\leq AM^{-d}.$
For any set of edges $\gamma$, we define
$N(\gamma):=\sum_{e\in\gamma}I_{e,M},\quad N_{L,M}:=\max_{\gamma\in\Xi_{L}}N(\gamma),$
where for $L\geq 1$,
$\Xi_{L}:=\{\gamma:\gamma\text{ is a set of edges in }\Lambda_{L};\,|\gamma|\leq L\}.$
Lemma 3.1.
[CN19, Lemma 2.6] Let $M\geq 1,a,A>0$ and $(I_{e,M})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be a collection of random variables satisfying (E1) and (E2). Then, there exists a positive constant $C=C(a,A,d)$ such that
  1. (i)
    For all $L\geq 1$,
    ${\mathbb{E}}[N_{L,M}]\leq CLq_{M}^{1/d}M^{d+1}.$
  2. (ii)
    If $t\geq CM^{d}\max\left(1,MLq_{M}^{1/d}\right)$, then
    $\mathbb{P}(N_{L,M}\geq t)<2^{d}\exp(-t/(16M)^{d}).$
Remark 3.2.
In [CN19, Lemma 2.6], although the result is stated for site-percolation, it also holds true for edge-percolation by similar arguments, see for example [CNN24, Lemma 2.6].
We aim at extending this result to general weight distributions. Let $a$ and $A$ be positive constants. Suppose that $(X_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ is a collection of non-negative random variables satisfying the following: for all integer $M\geq 1$
  • (P1)
    for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, the event $\{M-1\leq X_{e}<M\}$ is independent of $(X_{e^{\prime}})_{e^{\prime}\notin\Lambda_{aM}(e)}$,
  • (P2)
    there exists a function $\phi:[0,\infty)\to[0,\infty)$ such that $\phi(M)\leq AM^{-d}$ and
    $q_{M}=\sup_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}\mathbb{P}(M-1\leq X_{e}<M)\leq\phi(M).$
Lemma 3.3.
Let $X=(X_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be a family of random variables satisfying (P1) and (P2) and let $f:[0,\infty)\to[0,\infty)$ be a function satisfying
$B:=\sum_{M=1}^{\infty}\left(f_{*}(M)+f_{*}^{2}(M)+f_{*}^{4}(M)\right)M^{d+1}\phi(M)^{1/d}<\infty,\quad f_{*}(M):=\sup_{M-1\leq x<M}f(x).$ (H)
Then there exists a positive constant $C=C(a,A,B)$ such that the following holds.
  1. (i)
    For all $L\geq 1$,
    ${\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]\leq CL^{2}.$
  2. (ii)
    Let $\gamma$ be a random path starting from 00 in the same probability space of $X$. Then for all $L\geq 1$,
    ${\mathbb{E}}\left[\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]\leq CL^{2}+C\sum_{\ell\geq L}\ell^{2}({\mathbb{P}}(|\gamma|=\ell))^{1/2}.$
  3. (iii)
    Let $m\geq 1$ and $\gamma$ be a random set of edges of $\Lambda_{m}$ in the same probability space of $X$. Then for all $L\geq 1$,
    ${\mathbb{E}}\left[\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]\leq C(L+m)^{2}+C\sum_{\ell\geq L}(\ell+m)^{2}({\mathbb{P}}(|\gamma|=\ell))^{1/2}.$
  4. (iv)
    Let $L\geq 1$ and $K\in{\mathbb{N}}$ such that $LM\phi(M)^{1/d}\geq 1$ for all $1\leq M\leq K$. Assume further that $X_{e}\leq K$ for all $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$. There exist constants $C_{1},c_{1}>0$ such that for any random path $\gamma$ starting from 00 in the same probability space of $X$,
    $\mathbb{P}\left(\sum_{e\in\gamma}f(X_{e})\geq C_{1}L\right)\leq\sum_{M=1}^{K}\exp(-c_{1}LM\phi(M)^{1/d})+\mathbb{P}(|\gamma|\geq L).$
Proof.
We first prove (i). By Cauchy-Schwarz inequality,
${\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]={\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]\leq{\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}|\gamma|\sum_{e\in\gamma}f^{2}(X_{e})\right]\leq L{\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f^{2}(X_{e})\right],$ (3.1)
since $|\gamma|\leq L$ for all $\gamma\in\Xi_{L}$. For set of edges $\gamma$, we define
$A^{\gamma}_{M}:=\left|\{e\in\gamma:M-1<X_{e}\leq M\}\right|=\sum_{e\in\gamma}I_{e,M},$
where
$I_{e,M}:={\mathbb{I}}(M-1<X_{e}\leq M).$
Notice that,
$\sum_{e\in\gamma}f^{2}(X_{e})\leq\sum_{M\geq 1}f^{2}_{*}(M)A^{\gamma}_{M}+f^{2}(0)|\gamma|,$ (3.2)
and hence
${\mathbb{E}}\left[\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f^{2}(X_{e})\right]$ $\leq{\mathbb{E}}\Bigg{[}\sum_{M\geq 1}f_{*}^{2}(M)\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}I_{e,M}\Bigg{]}+f^{2}(0)L$
$=\sum_{M\geq 1}f_{*}^{2}(M){\mathbb{E}}\left[N_{L,M}\right]+f^{2}(0)L,$ (3.3)
where
$N_{L,M}=\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}I_{e,M}.$
By (P1) and (P2), the conditions (E1) and (E2) are satisfied for all $M\geq 1$. Now using Lemma 3.1 (i), we obtain that for all $M\geq 1$,
${\mathbb{E}}[N_{L,M}]={\mathcal{O}}(1)LM^{d+1}\phi(M)^{1/d}.$ (3.4)
This together with (3) implies that
${\mathbb{E}}\Bigg{[}\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f^{2}(X_{e})\Bigg{]}$ $={\mathcal{O}}(1)L\sum_{M\geq 1}f_{*}^{2}(M)M^{d+1}(\phi(M))^{1/d}={\mathcal{O}}(L),$
using (H). Finally, combining the above estimate with (3.1), we obtain (i). For (ii), we decompose
${\mathbb{E}}\left[\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]$ $={\mathbb{E}}\left[\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}{\mathbb{I}}(|\gamma|<L)\right]+{\mathbb{E}}\left[\left(\sum_{e\in\gamma}f(X_{e})\right)^{2}{\mathbb{I}}(|\gamma|\geq L)\right]$
$\leq{\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{L}}\sum_{e\in\gamma}f(X_{e})\right)^{2}\right]+\sum_{\ell=L}^{\infty}{\mathbb{E}}\left[|\gamma|\sum_{e\in\gamma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right]$
$\leq{\mathcal{O}}(L^{2})+\sum_{\ell=L}^{\infty}\ell{\mathbb{E}}\left[\max_{\gamma\in\Xi_{\ell}}\sum_{e\in\gamma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right],$ (3.5)
by using (i). Moreover, by Cauchy-Schwarz inequality and (i),
${\mathbb{E}}\left[\max_{\gamma\in\Xi_{\ell}}\sum_{e\in\gamma}f^{2}(X_{e}){\mathbb{I}}(|\gamma|=\ell)\right]$ $\leq{\mathbb{E}}\left[\left(\max_{\gamma\in\Xi_{\ell}}\sum_{e\in\gamma}f^{2}(X_{e})\right)^{2}\right]^{1/2}{\mathbb{E}}\left[{\mathbb{I}}(|\gamma|=\ell)\right]^{1/2}$
$\leq{\mathcal{O}}(\ell)(\mathbb{P}(|\gamma|=\ell))^{1/2}.$
Combining the last two displays yields (ii). We can easily prove (iii) by using the same arguments as for (ii) and the fact that if $|\gamma|\leq t$ then $\gamma\in\Xi_{t+m}$ for all $t\geq 1$.
Finally, we show (iv). Using $X_{e}\leq K$ and a similar estimate as in (3.2), we have for any $C_{1}\geq 2f(0)$,
$\mathbb{P}\left(\sum_{e\in\gamma}f(X_{e})\geq C_{1}L\right)$ $\leq\mathbb{P}\left(\sum_{e\in\gamma}f(X_{e})\geq C_{1}L,|\gamma|\leq L\right)+\mathbb{P}\left(|\gamma|\geq L\right)$
$\leq\mathbb{P}\left(\sum_{M=1}^{K}f_{*}(M)N_{L,M}\geq C_{1}L/2\right)+\mathbb{P}\left(|\gamma|\geq L\right).$ (3.6)
Furthermore, the conditions (E1) and (E2) are satisfied for all $M\geq 1$. Let $C$ be the constant as in Lemma 3.1, and set
$C_{1}:=2f(0)+2C\sum_{M=1}^{\infty}f_{*}(M)(\phi(M))^{1/d}M^{d+1}.$
Note that $C_{1}\in(0,\infty)$ by (H). Using Lemma 3.1 (ii),
$\mathbb{P}\left(\sum_{M=1}^{K}f_{*}(M)N_{L,M}\geq C_{1}L/2\right)$ $\leq\mathbb{P}\left(\sum_{M=1}^{K}f_{*}(M)N_{L,M}\geq\sum_{M=1}^{K}Cf_{*}(M)\phi(M)^{1/d}LM^{d+1}\right)$
$\leq\sum_{M=1}^{K}\mathbb{P}\left(N_{L,M}\geq C\phi(M)^{1/d}LM^{d+1}\right)$
$\leq\sum_{M=1}^{K}\exp(-c_{1}LM\phi(M)^{1/d}),$
with $c_{1}=c_{1}(d)$ a positive constant. Combining this with (3), we derive (iv). ∎
Finally, we shall apply Lemma 3.3 to the sequence of truncated effective radii defined in (2.4): $\hat{R}_{e}=\min\{C_{*}R_{e},\log^{2}n\}$ for $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$.
Corollary 3.4.
There exists a positive constant $C$ such that the following holds for all integers $n,m$ sufficiently large.
  1. (i)
    Let $\gamma$ be a random path in the same probability space of $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$, starting from 00 and satisfying ${\mathbb{P}}(|\gamma|=\ell)\leq C\ell^{-7}$ for all integer $\ell\geq Cn$. Then
    ${\mathbb{E}}\Big{[}\Big{(}\sum_{e\in\gamma}\hat{R}_{e}\Big{)}^{2}\Big{]}\leq{\mathbb{E}}\Big{[}\Big{(}\sum_{e\in\gamma}\hat{R}^{2}_{e}\Big{)}^{2}\Big{]}\leq Cn^{2}.$
  2. (ii)
    Let $\gamma$ be a random set of edges of $\Lambda_{m}$ for some integer $m\geq 1$ in the same probability space of $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ satisfying ${\mathbb{P}}(|\gamma|=\ell)\leq C\ell^{-7}$ for all integer $\ell\geq Cm$. Then
    ${\mathbb{E}}\Big{[}\sum_{e\in\gamma}\hat{R}_{e}^{2}\Big{]}\leq Cm.$
  3. (iii)
    For any $c>0$, there exists $C_{1}=C_{1}(c)$ such that for all $\gamma$, a random path starting from 00 in the same probability space of $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$,
    $\mathbb{P}\left(\sum_{e\in\gamma}\hat{R}_{e}^{2}\geq C_{1}n\right)\leq C_{1}\exp\left(-\frac{n}{C_{1}\log^{2(d+10)}n}\right)+\mathbb{P}(|\gamma|\geq cn).$
Proof.
Let $f(x)=x^{2}$ and $\phi(x)=x^{-d^{2}+11d}$, so (H) is verified. By Remark 2.3 and Proposition 2.4, the radii $(\hat{R}_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ satisfy (P1) and (P2). Thus, the parts (i) and (ii) follow from Lemma 3.3 (ii) and (iii), respectively. Finally, applying Lemma 3.3 (iv) to $L=cn,K=\lceil\log^{2}n\rceil$, we obtain (iii). ∎
4. Subdiffusive concentration of ${\mathrm{T}}_{n}$
The proof strategy for Theorem 1.3 relies on establishing a connection between bounds on $\operatorname{Var}(e^{\lambda{\mathrm{T}}_{n}/2})$ and exponential concentration (see [BR08, Lemma 4.1]). To achieve the required variance bound, we employ the Falik-Samorodnitsky inequality (Lemma 4.1) to a martingale decomposition of the random variable $e^{\lambda{\mathrm{F}}_{m}}$, where ${\mathrm{F}}_{m}$ represents an averaged version of the passage time. This approach was first introduced by Benaïm and Rossignol in [BR08] and subsequently utilized by Damron, Hanson, and Sosoe in [DHS14]. Finally, we estimate the tails of the true passage time ${\mathrm{T}}_{n}$ based on those of ${\mathrm{F}}_{m}$.
4.1. Variance bound via entropy inequality
Let us enumerate the edges ${{\mathcal{E}}({\mathbb{Z}}^{d})}$ as $\{e_{1},e_{2},\ldots\}$ and let $a,b\in{\mathbb{R}}\cup\{+\infty\}$. Assume that $(t_{e_{i}})_{i\geq 1}$ are i.i.d. random variables with the same distribution as
$\zeta=p\delta_{a}+(1-p)\delta_{b}.$
Let $g:\{a,b\}^{{\mathcal{E}}({\mathbb{Z}}^{d})}\to{\mathbb{R}}$ be a function of $(t_{e_{i}})_{i\geq 1}$. Fix $\lambda\in{\mathbb{R}}$ and define
$G:=G_{\lambda}:=e^{\lambda g}.$
We write
$G=G(t_{e_{i}},t_{e_{i}^{c}})$
to emphasize $G$ is the function of the random variables $t_{e_{i}}$ and $t_{e_{i}^{c}}=(t_{e_{j}})_{j\neq i}$. We define the natural filtration of these random variables as
${\mathcal{F}}_{0}:=\{\varnothing,\Omega\},\quad{\mathcal{F}}_{i}:=\sigma(t_{e_{1}},\ldots,t_{e_{i}}),$
for each $i\geq 1$. Now we consider the martingale increments
$\forall i\geq 1,\quad\Delta_{i}:={\mathbb{E}}[G\mid{\mathcal{F}}_{i}]-{\mathbb{E}}[G\mid{\mathcal{F}}_{i-1}]={\mathbb{E}}[G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})\mid{\mathcal{F}}_{i-1}],$
where $t^{\prime}_{e_{i}}$ is an independent copy of $t_{e_{i}}$. We will bound the variance of $G$ based on the following entropy inequality by Falik and Samorodnitsky [FS07, Lemma 2.3].
Lemma 4.1 (Falik-Samorodnitsky).
If ${\mathbb{E}}[G^{2}]<\infty$ then
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]\geq\operatorname{Var}[G]\log\dfrac{\operatorname{Var}[G]}{\sum_{i=1}^{\infty}({\mathbb{E}}[|\Delta_{i}|])^{2}},$ (4.1)
where ${\operatorname{Ent}}$ denotes the entropy operator: for any non-negative random variable $X$ with finite moment,
${\operatorname{Ent}}[X]:={\mathbb{E}}\left[X\log\dfrac{X}{{\mathbb{E}}[X]}\right].$
The following estimate on the total entropy is derived from the tensorization property of entropy and the log-Sobolev inequality for the Bernoulli distribution.
Lemma 4.2.
Assume that ${\mathbb{E}}[G^{4}]<\infty$. Then, there exists a positive constant $C$ depending on $p$ such that
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]\leq C\sum_{i=1}^{\infty}{\mathbb{E}}[(G(b,t_{e_{i}^{c}})-G(a,t_{e_{i}^{c}}))^{2}].$ (4.2)
4.2. Proof of Theorem 1.3
Instead of directly showing the subdiffusive concentration of ${{\mathrm{T}}_{n}}$, we will employ a strategy inspired by Benjamini, Kalai, and Schramm in [BKS03], known as the BKS trick. This approach involves proving the subdiffusive concentration for a geometric average of passage time, a concept introduced in [BKS03] and later used in [AZ13, DHS14, DHS15, BDG20]. It is expected that the majority of edges in the lattice have a low probability of lying in the geodesic of ${\mathrm{T}}_{n}$, meaning they have a small influence. However, this does not hold for edges very close to the origin, and the BKS trick provides a way to circumvent this issue.
First of all, let us define a spatial average of the first passage time,
${{\mathrm{F}}_{m}}:=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathrm{T}}_{z},$ (4.3)
where
$\forall z\in{\mathbb{Z}}^{d},\quad{\mathrm{T}}_{z}:={\mathrm{T}}(z,z+n{\mathbf{e}}_{1}),\qquad m:=\lfloor n^{1/4}\rfloor.$
To prove Theorem 1.3, it suffices to show the following variance bound.
Theorem 4.3.
There exists a constant $c>0$ such that
$\forall\,|\lambda|<\dfrac{1}{\sqrt{K}},\quad\operatorname{Var}\left[e^{\lambda{{\mathrm{F}}_{m}}}\right]\leq K\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]<\infty,$ (4.4)
where $K=\dfrac{cn}{\log n}$.
The following result is a direct consequence of Theorem 4.3. We refer the reader to [BR08, Lemma 4.1] for a proof.
Corollary 4.4.
There exists a positive constant $c$ such that
$\mathbb{P}\left(\left|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]\right|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa)\quad\forall\kappa\geq 0.$ (4.5)
Next, we prepare a simple large deviation estimate for the first passage time, which will be used to compare ${\mathrm{T}}_{n}$ and ${\mathrm{F}}_{m}$.
Lemma 4.5.
There exist positive constants $C,c$ such that for all $x,y\in{\mathbb{Z}}^{d}$ and $t\geq C\|x-y\|_{\infty}$,
$\mathbb{P}({\mathrm{T}}(x,y)\geq t)\leq c^{-1}\exp(-ct/\log^{2}n).$ (4.6)
Proof.
Observe that ${\mathrm{T}}(u,v)\leq\log^{2}n\|u-v\|_{1}$ for all $u,v\in{\mathbb{Z}}^{d}$. Therefore, by the triangle inequality
$\mathbb{P}\left({\mathrm{T}}(x,y)\geq t\right)$ $\leq\mathbb{P}\left({\mathrm{T}}(x,x^{*})+{\mathrm{T}}(y,y^{*})+{\mathrm{T}}(x^{*},y^{*})\geq t\right)$
$\leq\mathbb{P}\left({\mathrm{D}}^{*}(x,y)\geq t/2\right)+\mathbb{P}\left(\|x-x^{*}\|_{1}\geq\tfrac{t}{4\log^{2}n}\right)+\mathbb{P}\left(\|y-y^{*}\|_{1}\geq\tfrac{t}{4\log^{2}n}\right),$
where we recall that $z^{*}$ is the closest point of $z$ in the infinite cluster ${\mathcal{C}}_{\infty}$. The last two terms are bounded by $c_{1}^{-1}\exp\left(\tfrac{-c_{1}t}{\log^{2}n}\right)$, for some positive constants $c_{1}$ using Lemma 2.8, whereas by Lemma 2.7, the first term is bounded by $c_{2}^{-1}\exp(-c_{2}t)$ when $t\geq C\|x-y\|_{\infty}$ with some $C,c_{2}>0$. Hence, the result follows. ∎
Proof of Theorem 1.3. Since ${\mathbb{E}}[{{\mathrm{F}}_{m}}]={\mathbb{E}}[{{\mathrm{T}}_{n}}]$,
$|{{\mathrm{T}}_{n}}-{\mathbb{E}}[{{\mathrm{T}}_{n}}]|=|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{T}}_{n}}]+{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|$ $=|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]+{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|$
$\leq|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]|+|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|.$ (4.7)
Thus, for all $M\geq 1$, using the union bound, we have
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{\mathbb{E}}[{{\mathrm{T}}_{n}}]|\geq 4M\right)\leq\mathbb{P}\left(|{{\mathrm{F}}_{m}}-{\mathbb{E}}[{{\mathrm{F}}_{m}}]|\geq 2M\right)+\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|\geq 2M\right).$ (4.8)
By subadditivity property,
$|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|=\Big{|}{{\mathrm{T}}_{n}}-\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathrm{T}}_{z}\Big{|}$ $\leq\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}|{\mathrm{T}}(0,n{\mathbf{e}}_{1})-{\mathrm{T}}(z,z+n{\mathbf{e}}_{1})|$
$\leq\dfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}(0,z)+{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)).$ (4.9)
Observe that if the event $\Big{\{}\tfrac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}(0,z)+{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z))\geq 2M\Big{\}}$ occurs,
$\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\text{ or }\max_{z\in\Lambda_{m}}{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)\geq M.$ (4.10)
Combining this with union bound, it yields that
$\mathbb{P}$ $\left(\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}(0,z)+{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z))\geq 2M\right)$
$\qquad\qquad\qquad\leq\mathbb{P}\left(\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\right)+\mathbb{P}\left(\max_{z\in\Lambda_{m}}{\mathrm{T}}(n{\mathbf{e}}_{1},n{\mathbf{e}}_{1}+z)\geq M\right)$
$\qquad\qquad\qquad=2\mathbb{P}\left(\max_{z\in\Lambda_{m}}{\mathrm{T}}(0,z)\geq M\right)\leq 2|\Lambda_{m}|\max_{z\in\Lambda_{m}}\mathbb{P}\left({\mathrm{T}}(0,z)\geq M\right),$ (4.11)
where for the equation we have used the translation invariance.
Let $M=\tfrac{1}{4}\sqrt{\frac{n}{\log n}}\kappa$. Since $m=o(M)$, Lemma 4.5 shows that
$\max_{z\in\Lambda_{m}}\mathbb{P}\left({\mathrm{T}}(0,z)\geq M\right)\leq c^{-1}\exp(-cM/\log^{2}n),$ (4.12)
for some constant $c>0$. Using this estimate, (4.2), and (4.2) yields
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-{{\mathrm{F}}_{m}}|\geq\dfrac{\kappa}{2}\sqrt{\tfrac{n}{\log n}}\right)\leq{\mathcal{O}}(m^{d})\exp\Big{(}-c\tfrac{\sqrt{n}}{4\sqrt{\log^{5}n}}\kappa\Big{)}.$
Combining this with Corollary 4.4 and (4.8), it follows that
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-\mathbb{E}[{{\mathrm{T}}_{n}}]|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa)\qquad\forall\kappa\geq 0,$ (4.13)
for some positive constant $c$. ∎
Proof of Theorem 4.3. According to the Lemma 4.1, the variance bound is induced by two crucial factors: the estimate of total influence and total entropy. These keys are presented in the following results.
Proposition 4.6.
There exists a positive constant $C$ such that
$\sum_{i=1}^{\infty}({\mathbb{E}}[|\Delta_{i}|])^{2}\leq C\lambda^{2}{\mathbb{E}}\big{[}e^{2\lambda{\mathrm{F}}_{m}}\big{]}n^{(9-d)/8}\quad\forall\,\lambda\in{\mathbb{R}}.$
Proposition 4.7.
There exists a positive constant $C$ such that
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq C\lambda^{2}n{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\quad\forall\,|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}.$ (4.14)
By Lemma 4.1, Propositions 4.6 and 4.7, we have
$\operatorname{Var}\big{[}e^{\lambda{\mathrm{F}}_{m}}\big{]}\leq C\left(\log\frac{\operatorname{Var}[e^{\lambda{\mathrm{F}}_{m}}]}{C\lambda^{2}n^{(9-d)/8}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]}\right)^{-1}\lambda^{2}n{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$ (4.15)
We can assume that
$\operatorname{Var}[e^{\lambda{\mathrm{F}}_{m}}]\geq C\lambda^{2}n^{15/16}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right],$ (4.16)
since otherwise there is nothing left to prove. By (4.15) and (4.16), there exists a constant $c>0$ such that for any $|\lambda|\leq\tfrac{1}{\log^{2(d+11)}n}$,
$\operatorname{Var}\big{[}e^{\lambda{\mathrm{F}}_{m}}\big{]}\leq c\lambda^{2}\dfrac{n}{\log n}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$
This concludes the proof of Theorem 4.3 by substituting $\lambda/2$ for $\lambda$. ∎
In the rest of Section 4, we prove Propositions 4.6 and 4.7 in subsections 4.3 and 4.4, respectively.
4.3. Bound on the total influence: Proof of Proposition 4.6
Proposition 4.6 is a direct consequence of the following lemma with the notice that $m=\lfloor n^{1/4}\rfloor$.
Lemma 4.8.
There exists a constant $C>0$ such that the following holds.
  • (i)
    $\sup_{i\geq 1}{\mathbb{E}}[|\Delta_{i}|]\leq C|\lambda|m^{(1-d)/2}\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\quad\forall\,\lambda\in{\mathbb{R}}.$ (4.17)
  • (ii)
    $\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]\leq C|\lambda|n\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\quad\forall\,\lambda\in{\mathbb{R}}.$ (4.18)
4.3.1. Proof of Lemma 4.8 (i)
Fix $i\geq 1$ and consider
$\Delta_{i}={\mathbb{E}}[G|{\mathcal{F}}_{i}]-{\mathbb{E}}[G|{\mathcal{F}}_{i-1}]={\mathbb{E}}\left[G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})|{\mathcal{F}}_{i-1}\right].$ (4.19)
We have
${\mathbb{E}}\left[|\Delta_{i}|\right]\leq{\mathbb{E}}\left[|G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})|\right]=2{\mathbb{E}}\left[\big{(}e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}-e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\big{)}_{+}\right],$
where $t^{\prime}_{e_{i}}$ is the independent copy of $t_{e_{i}}$. Furthermore, using the inequality that $(e^{\lambda a}-e^{\lambda b})_{+}\leq|\lambda|(e^{\lambda a}+e^{\lambda b})|a-b|$, we get
${\mathbb{E}}[|\Delta_{i}|]$ $\leq 2|\lambda|{\mathbb{E}}\left[\left(e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}+e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\right)\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\right]$
$=4|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\right]$
$\leq 8|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\right)\right],$ (4.20)
where for the last line we have used $\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\leq 2\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})\right)$. For each $z\in{\mathbb{Z}}^{d}$, let $\gamma_{z}$ be a geodesic of ${\mathrm{T}}_{z}(1,t_{e^{c}_{i}})$. By Proposition 2.7,
${\mathrm{T}}_{z}(\log^{2}n,t_{e^{c}_{i}})-{\mathrm{T}}_{z}(1,t_{e^{c}_{i}})\leq(\log^{2}n{\mathbb{I}}({\mathcal{U}}_{z,e_{i}})+\hat{R}_{e_{i}}){\mathbb{I}}(e_{i}\in\gamma_{z}),$
where recall that
$\mathcal{U}_{z,e_{i}}=\{R_{e_{i}}\geq r_{z,e_{i}}\},\quad r_{z,e_{i}}=\tfrac{1}{3}\|e_{i}-z\|_{\infty}\wedge\|e_{i}-(z+n{\mathbf{e}}_{1})\|_{\infty},\quad\hat{R}_{e_{i}}=\min\{C_{*}R_{e_{i}},\log^{2}n\}.$
Therefore,
${{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}_{z}(\log^{2}n,t_{e^{c}_{i}})-{\mathrm{T}}_{z}(1,t_{e^{c}_{i}}))$
$\leq\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\big{(}\log^{2}n{\mathbb{I}}({\mathcal{U}}_{z,e_{i}})+\hat{R}_{e_{i}}\big{)}{\mathbb{I}}(e_{i}\in\gamma_{z}).$
Observe that if the event ${\mathcal{U}}_{z,e_{i}}\cap\{r_{z,e_{i}}\geq\log^{3}n\}$ occurs, then $C_{*}R_{e_{i}}\geq\log^{2}n$ and so $\hat{R}_{e_{i}}=\log^{2}n$. Therefore, the above estimate implies that
$0\leq{{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})\leq A_{i},$ (4.21)
where
$A_{i}$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\big{(}2\hat{R}_{e_{i}}+\log^{2}n{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\big{)}{\mathbb{I}}(e_{i}\in\gamma_{z}).$ (4.22)
Combining this with (4.3.1) and Cauchy-Schwarz inequality yields
${\mathbb{E}}[|\Delta_{i}|]$ $\leq 8|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}}A_{i}\right]$ (4.23)
$\leq 8|\lambda|{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]^{1/2}{\mathbb{E}}\left[A_{i}^{2}\right]^{1/2}.$ (4.24)
Here for the first line, we remark that ${{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})={{\mathrm{F}}_{m}}$.
Next, we will estimate ${\mathbb{E}}[A_{i}^{2}]$. Notice that for all edges $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ and $\Lambda\subset{\mathbb{Z}}^{d}$,
|$\left|\{z\in\Lambda:r_{z,e}\leq t\}\right|\leq\left|\{z\in{\mathbb{Z}}^{d}:r_{z,e}\leq t\}\right|={\mathcal{O}}(t^{d}).$ (4.25)
Therefore,
$\frac{1}{\Lambda_{m}}\sum_{z\in\Lambda_{m}}{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\leq\frac{{\mathcal{O}}(\log^{3d}n)}{|\Lambda_{m}|}={\mathcal{O}}(m^{1-d}),$ (4.26)
since $m=\lfloor n^{1/4}\rfloor$. Thus, by Cauchy-Schwarz inequality,
$A_{i}^{2}$ $\leq\frac{8}{\Lambda_{m}}\sum_{z\in\Lambda_{m}}\hat{R}_{e_{i}}^{2}{\mathbb{I}}(e_{i}\in\gamma_{z})+{\mathcal{O}}(m^{2-2d}).$ (4.27)
Combining this estimate with the translation invariance, we have
${\mathbb{E}}[A_{i}^{2}]$ $\leq\frac{8}{|\Lambda_{m}|}{\mathbb{E}}\left[\sum_{z\in\Lambda_{m}}\hat{R}_{e_{i}-z}^{2}{\mathbb{I}}(e_{i}-z\in\gamma_{0})\right]+{\mathcal{O}}(m^{2-2d})$
$=\frac{8}{|\Lambda_{m}|}{\mathbb{E}}\left[\sum_{e\in\gamma}\hat{R}_{e}^{2}\right]+{\mathcal{O}}(m^{2-2d}),$ (4.28)
where
$\gamma:=\gamma_{0}\cap\{e_{i}-\Lambda_{m}\},\quad\{e_{i}-\Lambda_{m}\}:=\{e^{\prime}=(x_{e_{i}}-z,y_{e_{i}}-z):z\in\Lambda_{m}\}.$
Observe that if $|\gamma|\geq\ell$ then we can find $x,y$ in $V(e_{i},m)$, the vertex set of $e_{i}-\Lambda_{m}$, such that ${\mathrm{T}}(x,y)\geq\ell$. Therefore, using the union bound and Lemma 4.5 we have for all $\ell\geq Cm$ with $C$ a sufficiently large constant,
$\mathbb{P}(|\gamma|\geq\ell)\leq\mathbb{P}\left(\exists\,x,y\in V(e_{i},m):{\mathrm{T}}(x,y)\geq\ell\right)$ $\leq{\mathcal{O}}(1)m^{2d}\max_{x,y\in V(e_{i},m)}\mathbb{P}\left({\mathrm{T}}(x,y)\geq\ell\right)$
$\leq{\mathcal{O}}(1)\exp\left(-\frac{c\ell}{\log^{2}n}\right),$ (4.29)
with some positive constant $c>0$. Here, notice that to apply Lemma 4.5, we have used $\|x-y\|_{\infty}\leq 2m$ for all $x,y\in V(e_{i},m)$. The above estimate verifies the condition in Corollary 3.4 (ii) and thus
${\mathbb{E}}\Big{[}\sum_{e\in\gamma}\hat{R}_{e}^{2}\Big{]}={\mathcal{O}}(m).$ (4.30)
Combining (4.3.1) and (4.30) yields that for all $i\geq 1$,
${\mathbb{E}}[A_{i}^{2}]={\mathcal{O}}(m^{1-d}).$ (4.31)
Finally, we conclude from (4.24) and (4.31) that
$\sup_{i\geq 1}{\mathbb{E}}[|\Delta_{i}|]\leq{\mathcal{O}}(1)|\lambda|\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}m^{(1-d)/2},$ (4.32)
and the result follows. ∎
4.3.2. Proof of Lemma 4.8 (ii)
Using (4.23) and Cauchy-Schwarz inequality, we obtain that
$\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]$ $\leq 8|\lambda|{\mathbb{E}}\Big{[}e^{\lambda{{\mathrm{F}}_{m}}}\sum_{i=1}^{\infty}A_{i}\Big{]}\leq 8|\lambda|\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\Big{(}{\mathbb{E}}\Big{[}\Big{(}\sum_{i=1}^{\infty}A_{i}\Big{)}^{2}\Big{]}\Big{)}^{1/2},$ (4.33)
where $A_{i}$ is defined as in (4.22). Notice that
$\sum_{i=1}^{\infty}A_{i}$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{i=1}^{\infty}\big{(}2\hat{R}_{e_{i}}+\log^{2}n{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\big{)}{\mathbb{I}}(e_{i}\in\gamma_{z})$
$=\frac{2}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{e\in\gamma_{z}}\hat{R}_{e}+\frac{\log^{2}n}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{e\in\gamma_{z}}{\mathbb{I}}(r_{z,e}\leq\log^{3}n)$
$=\frac{2}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{e\in\gamma_{z}}\hat{R}_{e_{i}}+{\mathcal{O}}(\log^{3d+2}n),$
by using (4.25). Therefore, thanks to Cauchy-Schwarz inequality,
$\Big{(}\sum_{i=1}^{\infty}A_{i}\Big{)}^{2}\leq\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\Big{(}\sum_{e\in\gamma_{z}}\hat{R}_{e}\Big{)}^{2}+{\mathcal{O}}(\log^{6d+4}n).$ (4.34)
It follows from Lemma 4.5 that
$\mathbb{P}(|\gamma_{z}|\geq Cn)\leq c^{-1}\exp(-cn/\log^{2}n).$
Then applying Corollary 3.4 (i) gives
${\mathbb{E}}\left[\left(\sum_{e\in{\gamma}_{z}}\hat{R}_{e}\right)^{2}\right]={\mathcal{O}}(n^{2}).$ (4.35)
Combining (4.34) with (4.35) yields
${\mathbb{E}}\left[\left(\sum_{i=1}^{\infty}A_{i}\right)^{2}\right]={\mathcal{O}}(n^{2}),$ (4.36)
which together with (4.33) implies that
$\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]$ $\leq{\mathcal{O}}(1)|\lambda|n\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}.$
4.4. Entropy bound: Proof proposition 4.7
Using Lemma 4.2, the total entropy is bounded by
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq C\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}(G(\log^{2}n,t_{e_{i}^{c}})-G(1,t_{e_{i}^{c}}))^{2}\Big{]}=C\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{\lambda{\mathrm{F}}_{m}(\log^{2}n,t_{e_{i}^{c}})}-e^{\lambda{\mathrm{F}}_{m}(1,t_{e_{i}^{c}})}\Big{)}^{2}\Big{]}$
$\leq 2C|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{2\lambda{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})}+e^{2\lambda{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})}\Big{)}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$ (4.37)
Here for the last line we used the inequality that $|e^{\lambda a}-e^{\lambda b}|\leq|\lambda|(e^{\lambda a}+e^{\lambda b})(a-b)$ for all $a\geq b$. We remark that ${{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})$ and ${{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})$ are independent of $t_{e_{i}}$, and hence
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}$
$=\frac{1}{1-p}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}{\mathbb{I}}(t_{e_{i}}=\log^{2}n)\Big{]}$
$\leq\frac{1}{1-p}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$
Similarly,
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})}$ $\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}\leq\frac{1}{p}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$
Combining these inequalities with (4.4) and (4.21), we get
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq{\mathcal{O}}(1)\lambda^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}$
$\leq{\mathcal{O}}(1)\lambda^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}A_{i}^{2}\big{]}.$ (4.38)
By Cauchy-Schwarz inequality and (4.25),
$\sum_{i=1}^{\infty}A^{2}_{i}$ $\leq\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{i=1}^{\infty}\big{(}8\hat{R}^{2}_{e_{i}}+2\log^{4}n{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\big{)}{\mathbb{I}}(e_{i}\in\gamma_{z})$
$=\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{e\in\gamma_{z}}\hat{R}^{2}_{e}+\frac{2\log^{4}n}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\sum_{e\in\gamma_{z}}{\mathbb{I}}(r_{z,e}\leq\log^{3}n)$
$=\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}Y_{z}+{\mathcal{O}}(\log^{3d+4}n),$ (4.39)
where
$Y_{z}:=\sum_{e\in\gamma_{z}}\hat{R}^{2}_{e}.$
Combining the last two estimates, we obtain
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq\frac{{\mathcal{O}}(1)\lambda^{2}}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}\big{]}+{\mathcal{O}}(\log^{3d+4}n)\lambda^{2}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$ (4.40)
By Lemma 4.5 and the union bound, there exist positive constants $C^{\prime},c$ such that for all $z\in{\mathbb{Z}}^{d}$ and $t\geq C^{\prime}n$,
$\max\{\mathbb{P}({\mathrm{T}}_{z}\geq t),\mathbb{P}({\mathrm{F}}_{m}\geq t)\}\leq c^{-1}\exp\left(-\frac{ct}{\log^{2}n}\right).$ (4.41)
Particularly,
$\mathbb{P}(|\gamma_{z}|\geq t)\leq\mathbb{P}({\mathrm{T}}_{z}\geq t)\leq c^{-1}\exp\left(-\frac{ct}{\log^{2}n}\right).$ (4.42)
Using this estimate and Corollary 3.4 (iii), we obtain that
$\mathbb{P}(Y_{z}\geq Cn)\leq C\exp\left(-\frac{n}{C\log^{2(d+10)}n}\right),$ (4.43)
for some positive constant $C$. Moreover,
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}\Big{]}$ $\leq Cn{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\Big{]}+{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}{\mathbb{I}}(Y_{z}\geq Cn)\Big{]}$
$\leq Cn{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\Big{]}+({\mathbb{E}}[e^{8\lambda{{\mathrm{F}}_{m}}}])^{1/4}(\mathbb{P}(Y_{z}\geq Cn))^{1/4}({\mathbb{E}}[Y_{z}^{2}])^{1/2},$ (4.44)
thanks to Cauchy-Schwarz inequality. By (4.42) and Corollary 3.4 (i),
${\mathbb{E}}[Y_{z}^{2}]\leq{\mathcal{O}}(n^{2}).$ (4.45)
It follows from (4.41) that for all $\lambda\leq\frac{1}{\log^{2(d+11)}n}$,
${\mathbb{E}}[e^{8\lambda{{\mathrm{F}}_{m}}}]\leq 2\exp\left(\frac{8C^{\prime}n}{\log^{2(d+11)}n}\right),$
where $C^{\prime}$ is the constant in (4.41). Putting this estimate, (4.43), (4.4), and (4.45) together, we have for all $\lambda\leq\frac{1}{\log^{2(d+11)}n}$,
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}{\mathbb{I}}(Y_{z}\geq Cn)\Big{]}\leq C\exp\left(-\frac{n}{2C\log^{2(d+10)}n}\right).$ (4.46)
Thanks to (4.41) again, we have for all $-\frac{1}{\log^{2(d+11)}n}\leq\lambda<0$,
${\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\geq{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}{\mathbb{I}}({{\mathrm{F}}_{m}}\leq C^{\prime}n)\big{]}\geq\frac{1}{2}\exp\left(-\frac{2C^{\prime}n}{\log^{2(d+11)}n}\right),$
with $C^{\prime}$ being the constant defined as in (4.41). Combining the last two display equations with (4.4) yields for all $|\lambda|\leq\frac{1}{\log^{2(d+11)}n}$,
${\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}\big{]}$ $\leq Cn{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}+C\exp\left(-\frac{n}{2C\log^{2(d+10)}n}\right)\leq 2Cn{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]},$
which together with (4.40) implies the desired result. ∎
5. Comparison of the graph distance and the first truncated passage time
In order to prove Theorem 1.4, our aim is to upper bound the difference between ${\mathrm{D}}^{*}_{n}$ and ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$ by the total size of a collection of bypasses avoiding all closed ($\log^{2}n$-weight) edges along $\gamma$-a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. More precisely, we introduce an inductive process that reveals a set of closed edges $\Gamma\subset{\mathrm{clo}}(\gamma)$ that are separable, as described by property (ii) below. Consequently, the discrepancy can be bounded via the total sum of effective radii associated with edges in $\Gamma$. Recall that $C_{*}$ be the constant and $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be the effective radii defined as in Section 2.
Proposition 5.1.
Let $\gamma$ be a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. Then on the event $\{\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\,\forall e\in\mathrm{clo}(\gamma)\}$, there exists a subset $\Gamma\subset{\mathrm{clo}}(\gamma)$ such that
  1. (i)
    for all $e\in\Gamma$, $R_{e}\geq\tfrac{\log^{2}n}{2C_{*}}$,
  2. (ii)
    for all $e\neq e^{\prime}\in\Gamma$,
    $\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\},$
  3. (iii)
    $|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\leq 2C_{*}\sum_{e\in\Gamma}R_{e}.$
Proof.
Let us first show that the following procedure (G) is well defined:
Input: A path $\eta\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ satisfying ${\mathrm{clo}}(\eta)\neq\varnothing$ and $\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\,\forall e\in\mathrm{clo}(\eta)\}$.
Output: An edge $e_{*}=e_{*}(\eta)\in{\mathrm{clo}}(\eta)$ and a path $\eta_{e_{*}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ such that
  • (${\rm a}$)
    $e_{*}\in\arg\max\{R_{e}:e\in{\mathrm{clo}}(\eta)\}$;
  • (${\rm b}$)
    $\eta_{e_{*}}\setminus\eta$ is open and $R_{e_{*}}\geq\log^{2}n/2C_{*}$;
  • (${\rm c}$)
    $\text{clo}(\eta_{e_{*}})\cap\Lambda_{R_{e_{*}}}({e_{*}})=\varnothing$;
  • (${\rm d}$)
    ${\mathrm{T}}(\eta_{e_{*}}\setminus\eta)=|\eta_{e_{*}}\setminus\eta|\leq 2C_{*}R_{e_{*}}$.
Indeed, we simply take $e_{*}\in{\mathrm{clo}}(\gamma)$ with the largest effective radius, using a rule to break ties when there are several maximizing edges. Applying Proposition 2.6 to $\eta\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ and $e_{*}$, we obtain $\eta_{e_{*}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$, a modified path of $\eta$, satisfying $({\rm c})$, $({\rm d})$ and that $\eta_{e_{*}}\setminus\eta$ is open, and $|\eta_{e_{*}}\setminus\eta|\leq 2C_{*}R_{e_{*}}$. Therefore,
${\mathrm{T}}(\eta)\leq{\mathrm{T}}(\eta_{e_{*}})\leq{\mathrm{T}}(\eta)+{\mathrm{T}}(\eta_{e_{*}}\setminus\eta)-t_{e_{*}}\leq{\mathrm{T}}(\eta)+2C_{*}R_{e_{*}}-\log^{2}n,$
which implies $R_{e_{*}}\geq\log^{2}n/2C_{*}$ and thus (b) follows.
Figure 4. Illustration for covering process of all closed edges in the original geodesic $\gamma$
We set $\eta^{0}=\gamma$ and iteratively apply the procedure (G) to obtain a sequence of edges $(e_{k})_{k=1}^{\ell}\subset{\mathrm{clo}}(\eta^{0})$ and paths $(\eta^{k})_{k=1}^{\ell}$ such that $e_{k}=e_{*}(\eta^{k-1})$, $\eta^{k}=(\eta^{k})_{e_{k}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ for $k=1,\ldots,\ell$, and that (a)-(d) hold for each pair of $(e_{k},\eta^{k})_{k=1}^{\ell}$ and more importantly $\eta^{\ell}$ is open (see Figure 4 for illustration). Remark that the iteration stops after some finite steps $\ell$, since $|{\mathrm{clo}}(\eta^{0})|=|{\mathrm{clo}}(\gamma)|$ is finite and $|{\mathrm{clo}}(\eta^{k})|<|{\mathrm{clo}}(\eta^{k-1})|$ using (b) and (c). Setting $\Gamma=\{e_{1},\ldots,e_{\ell}\}$, we can easily deduce (i)–(iii) from (a)–(d). Indeed, (i) follows from (b). By (a) and (b), the sequence $(R_{e_{k}})_{k=1}^{\ell}$ is decreasing and hence by (c)
$\forall 1\leq i<j\leq\ell,\quad\|e_{i}-e_{j}\|_{\infty}\geq R_{e_{i}}=\max\{R_{e_{i}},R_{e_{j}}\}.$
Since $\eta^{\ell}$ is an open path from $0^{*}$ to $(ne_{1})^{*}$, the property (${\rm d}$) shows
$0\leq{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq{\mathrm{T}}(\eta^{\ell})-{\mathrm{T}}(\eta^{0})\leq\sum_{i=1}^{\ell}{\mathrm{T}}(\eta^{i})-{\mathrm{T}}(\eta^{i-1})\leq\sum_{i=1}^{\ell}|\eta^{i}\setminus\eta^{i-1}|\leq 2C_{*}\sum_{e\in\Gamma}R_{e},$
which implies (iii). ∎
Proof of Theorem 1.4. It is enough to prove the following statement: there exists a positive constant $c$ such that for all $L\geq\log^{2}n,$
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\geq L\right)\leq c^{-1}\exp(-cL/\log L).$ (5.1)
Indeed, by (5.1) and Lemma 2.8,
$\mathbb{P}(|{\mathrm{D}}_{n}^{*}-{\mathrm{T}}_{n}|\geq L)$ $\leq\mathbb{P}(|{\mathrm{D}}_{n}^{*}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\geq L/2)+\mathbb{P}(\|0^{*}\|_{1}+\|(n{\mathbf{e}}_{1})^{*}-n{\mathbf{e}}_{1}\|_{1}\geq\tfrac{L}{2\log^{2}n})$
$\leq c^{-1}\exp(-cL/\log L)+2\mathbb{P}(\|0^{*}\|_{1}\geq\tfrac{L}{4\log^{2}n})$
$\leq c_{1}^{-1}\exp\left(-c_{1}\tfrac{L}{\log L+\log^{2}n}\right),$
where $c_{1}$ is a positive constant. The remainder of this section is devoted to proving (5.1). Let $\gamma_{n}$ be a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. By Lemma 2.8 and Lemma 2.10, there exist a positive constant $C$ such that
$\mathbb{P}({\mathrm{D}}^{*}_{n}\leq Cn/2;\|0-0^{*}\|_{\infty},\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\leq n/4)\geq 1-C\exp(-n/C).$ (5.2)
Remark further that ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq{\mathrm{D}}^{*}_{n}$, and if ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq k$ then $\gamma_{n}\subset\Lambda_{k}(0^{*})$, since $t_{e}\geq 1$ for all $e$. Hence, if the event in (5.2) holds then $\gamma_{n}\subset\Lambda_{Cn}$, and thus we have
$\mathbb{P}({\mathcal{E}})\geq 1-C\exp(-n/C),\qquad{\mathcal{E}}:=\{\gamma_{n}\subset\Lambda_{Cn}\}.$ (5.3)
We also define
:=${\mathcal{E}}_{*}:=\{\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\forall\,e\in\gamma_{n}\}.$
Using (5.2), (5.3), and Proposition 2.4, there are positive constants $C$ and $c$ such that
${\mathbb{P}}({\mathcal{E}}_{*}^{c})$ $\leq{\mathbb{P}}({\mathcal{E}}_{*}^{c}\cap{\mathcal{E}}\cap\{\|0-0^{*}\|_{\infty}+\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\leq n/2\})+{\mathbb{P}}({\mathcal{E}}^{c})$
+($\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\mathbb{P}(\{\|0-0^{*}\|_{\infty}+\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\geq n/2)$
$\leq{\mathbb{P}}(\exists\,e\in\Lambda_{Cn}:R_{e}\geq n/12)+3C\exp(-cn)\leq 4C\exp(-cn).$ (5.4)
Case 1: $L\geq Cn$. Using Lemma 2.10,
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\geq L\right)\leq\mathbb{P}\left({{\mathrm{D}}^{*}_{n}}\geq L\right)\leq\exp(-cL),$
and the result follows.
Case 2: $L<Cn$. Using (5.3), (5), Propositions 2.4 and 5.1,
$\mathbb{P}$ $\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\geq L\right)$
[…]
where $c$ is a positive constant and
  • (a)
    $\frac{\log^{2}n}{2C_{*}}\leq R_{e}\leq L$ for all $e\in\Gamma_{n}$,
  • (b)
    $\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}$, for all $e,e^{\prime}\in\Gamma_{n}$,
  • (c)
    $\sum_{e\in\Gamma_{n}}R_{e}\geq L/2C_{*}$.
In order to estimate the last term of (5), we set $M_{0}=\lceil\log^{3/2}n\rceil$, and $M_{q}=M_{0}2^{q}$ for $1\leq q\leq a_{n}$ with $a_{n}=\lceil\log_{2}L-\log_{2}\log n\rceil$. Remark that $\log^{3/2}n\leq R_{e}\leq L\leq M_{a_{n}}$ for all $e\in\Gamma_{n}$ and thus,
$\sum_{e\in\Gamma_{n}}R_{e}\leq\sum_{q=0}^{a_{n}}2M_{q}N_{q},$
where for each $1\leq q\leq a_{n}$,
$N_{q}:=|\{e\in\Gamma_{n}:R_{e}\in[M_{q},2M_{q}]\}|.$
Therefore, if (c) occurs then there exists $0\leq q\leq a_{n}$ such that $N_{q}\geq B_{q}$ where for $1\leq q\leq a_{n}$,
$B_{q}:=\Big{\lfloor}\dfrac{L}{4C_{*}M_{q}\log_{2}L}\Big{\rfloor}.$
Hence, it follows from the union bound that
$\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a)--(c)}\right)\leq\sum_{q=0}^{a_{n}}\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a), (b)};\,N_{q}\geq B_{q}\right).$ (5.6)
Moreover, by (b), if $R_{e}$ and $R_{e^{\prime}}$ are in $[M_{q},2M_{q}]$ for some $e,e^{\prime}\in\Gamma_{n}$, then
$\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}\geq M_{q}.$ (5.7)
Therefore,
$\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a), (b)};\,N_{q}\geq B_{q}\right)\leq S_{q},$ (5.8)
where
$S_{q}:=$ ($\mathbb{P}\Big{(}\,\exists\,\{e_{1},\ldots,e_{B_{q}}\}\subset\Lambda_{Cn}:R_{e_{j}}\in[M_{q},2M_{q}]\,\forall\,j\in[B_{q}];\|e_{j}-e_{k}\|_{\infty}\geq M_{q}\,\forall j\neq k\in[B_{q}]\Big{)},$ (5.9)
here recall the notation $[a]=\{1,\ldots,a\}$. The following claim is straightforward.
Claim. Given $C$, there exists a constant $c=c(d,C)>0$ such that the following holds: for any $M\in{\mathbb{N}}$ and $E\subset{{\mathcal{E}}({\mathbb{Z}}^{d})}$ satisfying
$\|e-e^{\prime}\|_{\infty}\geq M,\quad\forall e,e^{\prime}\in E,$
we can find $E^{\prime}\subset E$ such that $|E^{\prime}|\geq c|E|$ and $\|e-e^{\prime}\|_{\infty}\geq 5CM$ for all $e,e^{\prime}\in E^{\prime}$.
By this claim, there exists a positive constant $c$ such that for any $1\leq q\leq a_{n}$ if the event in (5.9) occurs then we can find $E^{\prime}\subset\{e_{1},\ldots,e_{B_{q}}\}$ such that $|E^{\prime}|\geq\lfloor cB_{q}\rfloor$ and $\|e-e^{\prime}\|_{\infty}\geq 5C_{*}M_{q}$ for all $e,e^{\prime}\in E^{\prime}$. As a result, we have
$S_{q}$ ($\leq\mathbb{P}\left(\,\exists\,\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}:R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall j\in[C_{q}]\right)$
$\leq\sum_{\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}}\mathbb{P}\left(R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall j\in[C_{q}]\right),$ (5.10)
where $C_{q}=\lfloor cB_{q}\rfloor$ and
$\mathcal{T}_{q}=\{\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\subset\Lambda_{Cn}:\|e^{\prime}_{j}-e^{\prime}_{k}\|_{\infty}\geq 5C_{*}M_{q}\,\forall j\neq k\in[C_{q}]\}.$
We remark that the event $R_{e}\in[M_{q},2M_{q}]$ only depends on the state of edges in $\Lambda_{2C_{*}M_{q}}(e)$. Therefore, given $\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}$, the family of the events ($(\{R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\})_{i\in[C_{q}]}$ are independent. Hence,
$\mathbb{P}\left(R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall i\in[C_{q}]\right)$ $=\prod_{j=1}^{C_{q}}\mathbb{P}\left(R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\right)\leq\exp(-\alpha M_{q}C_{q}),$
where $\alpha$ is a positive constant. Here for the inequality we have used Proposition 2.4 with the remark that $M_{q}\leq M_{a_{n}}\leq n^{2}$. This estimate together with (5) yields for all $1\leq q\leq a_{n}$,
$S_{q}\leq|\mathcal{T}_{q}|\exp(-\alpha C_{q}M_{q})$ $\leq(4Cn)^{dC_{q}}\exp(-\alpha C_{q}M_{q})\leq\exp(-\alpha C_{q}M_{q}/2)\leq\exp\left(-\tfrac{c\alpha L}{16C_{*}\log_{2}L}\right).$
Combining the above estimate with (5.8), (5.6), and (5), we obtain (5.1). ∎
Acknowledgment. We would like to thank Shuta Nakajima for his valuable comments. V. Q. Nguyen is partially supported by the Master, PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF).